### v1

# Shaping the Safety Boundaries: Understanding and Defending Against Jailbreaks in Large Language Models

Abstract
Jailbreaking in Large Language Models (LLMs) is a major security concern as it can deceive LLMs to generate harmful text. Yet, there is still insufficient understanding of how jailbreaking works, which makes it hard to develop effective defense strategies. We aim to shed more light into this issue: we conduct a detailed large-scale analysis of seven different jailbreak methods and find that these disagreements stem from insufficient observation samples. In particular, we introduce safety boundary, and we find that jailbreaks shift harmful activations outside that safety boundary, where LLMs are less sensitive to harmful information. We also find that the low and the middle layers are critical in such shifts, while deeper layers have less impact. Leveraging on these insights, we propose a novel defense called Activation Boundary Defense (ABD), which adaptively constrains the activations within the safety boundary. We further use Bayesian optimization to selectively apply the defense method to the low and the middle layers. Our experiments on several benchmarks show that ABD achieves an average DSR of over 98% against various forms of jailbreak attacks, with less than 2% impact on the model’s general capabilities.
Shaping the Safety Boundaries: Understanding and Defending Against Jailbreaks in Large Language Models
Lang Gao1,2,  Xiangliang Zhang3, Preslav Nakov1, Xiuying Chen1222Corresponding author. 1MBZUAI  2Huazhong University of Science and Technology 3University of Notre Dame {Lang.Gao, Preslav.Nakov, Xiuying.Chen}@mbzuai.ac.ae xzhang33@nd.edu
1 Introduction
The widespread use of Large Language Models (LLMs) across various fields Kaddour et al. (2023) has raised concerns about the safety of the output and the robustness of the model. It has been shown that jailbreak attacks, which use crafted prompts to deceive LLMs into generating harmful content, can bypass LLM’s safety alignment Liu et al. (2024b); Zou et al. (2023), and a lot of research has focused on developing defense mechanisms and counter-prompts to mitigate such attacks Robey et al. (2023); Xie et al. (2024).
Figure 1: Projected activation space overview of Vicuna-7B-v1.3 across different layers. Harmful activations are observed to cluster together, and we define the surrounding boundary as the safety boundary. The attack arrow indicates that jailbreak prompts shift harmful activations into the benign space to evade safety checks.
Understanding the internal mechanism of how jailbreaks work is crucial for improving defenses and enhancing safety measures. However, studies explaining jailbreaks remain largely ambiguous and controversial Yu et al. (2024); Ball et al. (2024a); Lin et al. (2024). From an explainability perspective, debates revolve around which layers are most responsible for a successful attack: He et al. (2024) prioritize the low layers, Zhou et al. (2024); Shen et al. (2024) focus on the middle layers, while Li et al. (2024) highlight the deep layers. Another contradiction lies in whether activations gradually shift as layers deepen or abruptly transition in the middle layers Zhao et al. (2024a); Shen et al. (2024). Our experiments suggest that these inconsistencies may stem from the limited sample size of prompts, typically around 100, used in prior analyses. From a defense perspective, existing approaches often rely on additional training processes Zhao et al. (2024a) or probing a limited number of samples Shen et al. (2024), which affects the utility of LLMs.
Here we aim to better understand the mechanism of how jailbreaking works. In particular, we provide our explanation of jailbreaks based on a comprehensive analysis of over 30,000 samples, a significantly larger scale than previous studies. As an example, Figure 1 shows the two-dimensional projection of three types of prompts: benign prompts that contain no harmful information, harmful prompts that attempt to induce the LLM to generate harmful content but fail, and various jailbreak prompts that successfully induce harmful outputs, collected from different types of jailbreak attack methods. The projection is based on the last token representations across different layers, referred to as activations, as they capture the model’s overall understanding of the entire input sequence Radford et al. (2019); Zou et al. (2023). We can see that different prompts exhibit clustering effects, where the jailbreak activations are distinct from the activation space of harmful prompts. This indicates that in each layer, the LLM performs a rigorous safety check within a safety boundary, effectively containing most of the harmful activations within this controlled space. Jailbreak prompts, however, bypass these safety checks by shifting activations beyond the safety boundary into an alternative, unmonitored activation space. Notably, the extent of this shift varies across different layers of the LLM, with the low and the middle layers exhibiting the most pronounced effects. This suggests that these layers are critical to the success of jailbreaks.
This finding motivates a new defense mechanism that we propose, Activation Boundary Defense (ABD), which constrains jailbreak prompt activations within a safety boundary by applying penalties to activation values. Specifically, we impose minimal penalties within the boundary but sharply increase them beyond it, thereby preserving model utility. The constraint function considers various factors, including the defense layer and the defense dimension, which balance performances on defense and general tasks. We propose a novel adaptive objective that maximizes the Defense Success Rate (DSR) while minimizing the number of penalized layers. This objective is implemented with an adapted Bayesian Optimization (BO) method Jones et al. (1998); Mockus (1974), which iteratively proposes and refines selection suggestions.
Experiments on AdvBench Zou et al. (2023) show that ABD has good generalization, achieving an average DSR of over 98% against various forms of jailbreak attacks. In the general ability tasks, the performance drop of the model equipped with ABD does not exceed 2%, which is significantly lower compared to other defenses, such as 37% for Retokenization Jain et al. (2023). This demonstrates that our method has a minimal impact on the model’s overall abilities. In summary, our contributions are threefold:
  • We uncover a comprehensive activation distribution and introduce the safety boundary, resolving contradictions in prior work and highlighting the vulnerability of the low and the middle layers.
  • We propose a lightweight, extensible defense that penalizes only targeted samples and a few key layers, ensuring efficiency and precision.
  • Our experiments on benchmark datasets demonstrate that ABD achieves an average DSR of over 98% against various jailbreak attacks, with less than a 2% impact on the general capabilities.
2 Related Work
Jailbreak attacks on LLMs.
Jailbreak attacks craft prompts to induce harmful outputs from LLMs. Early approaches, like DAN Shen et al. (2023), rely on manually designed prompts. Later research proposed automatic methods. Optimization-based methods Zou et al. (2023); Liu et al. (2024b); Guo et al. (2024) design optimization strategies to iteratively edit and refine original harmful prompts, enhancing their stealthiness. Model-based methods Chao et al. (2023); Ding et al. (2024); Paulus et al. (2024) utilize an attacker LLM to autonomously query and refine jailbreak prompts. For rule-based jailbreak methods, Jiang et al. (2024) and Liu et al. (2024c) rewrite prompts using fixed rules that are sufficiently aggressive to deceive the model. In this work, we aim to analyze the mechanisms of these jailbreak methods to develop more effective defense strategies.
Defense against jailbreak attacks.
Unlike jailbreak methods, defense strategies enhance model safety and robustness by reformulating inputs to counter jailbreak prompts, such as backtranslation Wang et al. (2024), paraphrasing Jain et al. (2023), and reminding Xie et al. (2023). Some researchers use indicators to distinguish jailbreak prompts, e.g., classify input sequences based on perplexity and sequence length Alon and Kamfonas (2023), or classify jailbreak activations using linear classifiers to control whether to answer the query Zhao et al. (2024a). A new trend in defense involves directly manipulating the representations of the model Shen et al. (2024); Xu et al. (2024); Li et al. (2024); Liu et al. (2024a) and editing the model Zhao et al. (2024b), which is more efficient and cost-effective. Unlike these methods, ours avoids extra tokens or modules by directly constraining activations within a safety boundary, ensuring greater efficiency and simplicity.
Mechanistic interpretability of LLMs.
The growing concern about LLM safety has sparked increasing interest in interpreting LLM features in jailbreak prompts. For example, Ball et al. (2024b) identified a common mechanism whereby jailbreaks reduce the harmfulness perception in most LLMs. Similarly, Li et al. (2024) investigated patterns that trigger the model to recognize safety issues. Zhou et al. (2024) discovered that the vocabulary mappings of activations significantly changed when processing jailbreak inputs. Research efforts have also focused on proposing corresponding defense methods based on interpretability results. For instance, Zhao et al. (2024a) and Shen et al. (2024) model how jailbreak activations transfer between benign and harmful activation spaces as layers deepen. They designed adaptive defenses whose strength varies across different layers. However, the limited data in previous studies often led to controversy and ambiguity. In contrast, we leverage over 30,000 samples to better understand jailbreaks and design effective defenses.
3 Understanding Jailbreaks in LLMs
3.1 Controversy in Literature
What are the key layers?
There have been different opinions about which layers are most important, mainly due to varying perspectives. For example, He et al. (2024) argued that low layers are essential because they treat jailbreak activations as benign ones. Zhou et al. (2024) believe that low and middle layers are essential based on tokens generated from activations. Li et al. (2024) argued that deep layers are critical, as the concept they proposed, called safety pattern, has greater values in these layers. Among these works, the first work similarly focuses on activations, as does our paper.
Upon examining their experimental setup, we identified a major limitation: the small number of samples used in the analysis. As shown in Figure 2(a), before scaling the data, benign and harmful activations are linearly separable, with most jailbreak samples being misclassified as benign, which is consistent with the findings from He et al. (2024). However, after scaling the data, half of the jailbreak activations shift to align with harmful ones, and the separation becomes nonlinear. This highlights the risk of misleading from small-scale studies and underscores the importance of large-scale analysis.
Disputes on the jailbreak mechanism.
There are also different perspectives on the internal mechanism of jailbreak. Specifically, Zhao et al. (2024a) viewed jailbreak as a gradual process where activations transition from harmful to benign spaces from low to deep layers, while Shen et al. (2024) argued that jailbreak abruptly manifests in deeper layers, with activations positioned between harmful and benign spaces.
Figure 2: PCA visualization of the activations highlights how conclusions are drawn from tiny sample sizes, such as the side on which the jailbreak activation lies, may not hold with larger datasets. (a) Top: 100 samples per type (following He et al. (2024)); Bottom: Benign and harmful samples increased to 5000. (b) Top: 60 samples per type (similar to Zhao et al. (2024a)); Bottom: All samples increased to 500.
This work used no more than 100 samples per activation type, and we identify a similar issue to the one discussed earlier. Figure 2(b) shows that before scaling the data, the jailbreak activations are aligned to benign activations, which is consistent with findings from Zhao et al. (2024a); while after scaling the data, jailbreaking activations no longer align with either harmful or benign activations. When compared to Shen et al. (2024), similar conflicting perspectives are observed when scaling up the data. Further details can be found in Appendix C.1. These conflicting perspectives form the basis of the jailbreak mechanisms in these works; their conclusions — such as whether activations gradually shift or abruptly transition — are also questionable.
3.2 Experimental Settings
To address the above limitations, we conducted a large-scale analysis on a 300 times greater dataset.
Figure 3: (a) Impact of random activation shifts across layers. DSR (Defense Success Rate) decreases as shifting distance increases, regardless of the affected layers. (b) MVD (Most Vulnerable Distance) across layers. MVD increases as layers go deeper. (c) Inclusion ratio of jailbreaking activations in the harmful activation space. Without ABD, the ratio stays below 0.4 but rises to 1 when ABD is applied.
Dataset.
Our dataset consists of 32,507 samples in three categories: benign, harmful, and jailbreak. For benign samples, we used a subset from the Alpaca dataset Taori et al. (2023), as it has been carefully curated to ensure that only safe content is included. For harmful samples, we collected five datasets containing harmful queries from RedEval Bhardwaj and Poria (2023) and AdvBench Zou et al. (2023). These datasets encompass a wide variety of harmful prompts designed to elicit unsafe outputs. For jailbreak samples, we applied seven different jailbreak attack methods, including GCG-Individual and GCG-Universal Zou et al. (2023), AdvPrompter Paulus et al. (2024), COLD-Suffix Attack Guo et al. (2024), AutoDAN Liu et al. (2024b), PAIR Chao et al. (2023), and ArtPrompt Jiang et al. (2024), on AdvBench. Detailed statistics can be found in Appendix A.
Model.
We used Vicuna-7B-v1.3 Chiang et al. (2023), a 32-layer Transformer model, as the LLM to study. It is tuned upon a safety-aligned LLM and is able to identify certain harmful prompts. However, as demonstrated in previous work Chu et al. (2024), it can also be easily deceived and generate harmful information during jailbreak attacks. This makes it an excellent example for studying both attack mechanisms and defense strategies.
MDS projection.
Our projection was based on the activation, the last token vector of the input, as it captures the model’s overall understanding of the entire input sequence Radford et al. (2019); Zou et al. (2023). As discussed in §3.1, linear classification may be insufficient to model the boundary between different prompts. Therefore, we adopted Multi-Dimensional Scaling (MDS) Carroll and Arabie (1998) as our dimensionality reduction method, as MDS is more effective than PCA in handling large-scale and complexly distributed data Anowar et al. (2021). Moreover, it better captures the global distribution structure compared to t-SNE Van der Maaten and Hinton (2008).
3.3 Jailbreak Mechanisms Findings
3.3.1 Activation Distribution
We present the activation distributions of representative layers in Figure 1, with the full version available in Appendix B. Our key observations are as follows: (1) Harmful and benign activations overlap and are not linearly separable in most layers, especially in low and middle layers, as shown in Figure 1(a) and Figure 1(b), where benign and harmful activations are overlapping with no clear linear decision boundaries to separate them. This highlights the intricate interplay between harmful and benign activations, indicating that forcing jailbreak activations into a linear benign-vs-harmful classification is impractical. (2) Jailbreak activations significantly shift out of the harmful activation space in most layers, forming an independent region with low overlap with harmful activations. This pattern is consistent across various jailbreaks, suggesting that jailbreaks achieve deception by shifting activations. Such shift is evident in all layers, indicating that jailbreaks begin in low layers and continue to affect the model across all layers, rather than starting in middle or deep layers Zhou et al. (2024); Shen et al. (2024). Moreover, the shift is the most pronounced in the low and middle layers, implying their importance in jailbreaks.
3.3.2 Safety Boundary
Based on the observation of jailbreak and harmful activations, we propose the following hypothesis: A safety boundary exists that ensures sensitivity to harmful content within its activation space, but jailbreak prompts can bypass it, generating harmful responses. To verify the existence of the safety boundary, we conduct experiments and present two concepts described below.
We first explore Randomized Activation Shifting (RAS), an approximation of jailbreak attacks that introduces a random value shift to the original activation. Specifically, assuming the activation on the $l$-th layer is $a_{l}$, we shift it by a distance $r$ in a random direction $\hat{u}$:
$a_{l}\leftarrow a_{l}+r\cdot\hat{u}.$ (1)
Figure 3(a) shows the changes in Defense Success Rate (DSR) under different values of $r$ for four selected layers. The definition of this metric is provided in the Appendix. Initially, when $r=0$, the corresponding DSR is over 99%, as the harmful text has activations within the safety boundary. Then, as the distance $r$ increases, the DSR decreases for all layers, indicating that harmful activations become harder for the model to detect. Notably, regardless of the selected layer, a steep drop in DSR is observed at a specific range (e.g., $r=10-20$ on layer 0), while the decline is more moderate in other regions. This observation suggests the presence of a safety boundary.
For more accurate boundary detection, we provide the second visualization, which introduces a measurable concept called Most Vulnerable Distance (MVD) as an approximation of the boundary. MVD represents the specific distance at which the DSR experiences the steepest drop, indicating the point where the model’s ability to detect harmful activations becomes most vulnerable:
$\mathrm{MVD=\textstyle\arg\min_{r}}\textstyle\frac{\mathrm{d(DSR)}}{\mathrm{d}r},$ (2)
where $\frac{\mathrm{d(DSR)}}{\mathrm{d}r}$ denotes the rate of change of DSR with respect to $r$. As shown in Figure 3(b), MVD increases with the depth of the model’s layers. This implies that the safety boundary also expands in deeper layers, requiring more substantial shifts to compromise the model’s safety checks. The findings align closely with our intuitive observations in Figure 1. Figure 1(a) shows that in low layers, jailbreak activations are shifted in various directions, but all evidently surpass the safety boundary. In the middle layers presented in Figure 1(b), most jailbreak activations are beyond the safety boundary, and they are more aligned with benign activations. In the deep layers shown in Figure 1(c), jailbreak activations cluster within a region, which overlaps a little with the safety boundary.
3.3.3 Key Takeaways on the Jailbreak Process
Based on the above analysis, we summarize the jailbreak process as follows: (1) Jailbreak takes effect in the low layers, contrary to prior work Zhou et al. (2024), which suggested it begins in the middle layers. (2) As the attack progresses, jailbreak continues to operate, shifting the activation outside the safety boundary and causing the model to be deceived. (3) When comparing all layers, the low and the middle layers exhibit the strongest shift and smallest safety boundaries, demonstrating their importance in the jailbreak process.
4 Activation Boundary Defense
Based on the above findings, we propose a defense called Activation Boundary Defense (ABD), whose core idea is to confine the activations within the safety boundaries. The workflow of ABD is shown in Figure 4. ABD includes a penalty function and a Bayesian optimization process. The penalty constrains activations within safety boundaries, while Bayesian optimization iteratively adjusts the affected layers and the penalty parameters.
4.1 Penalty Function
Overall design of penalty.
Intuitively, an activation stays within the safety boundary if all its coordinates lie within a regular range; in contrast, outliers have at least one coordinate exceeding this range. Adjusting these outlier coordinates can guide the activations back within the boundary. Since directly measuring the safety boundaries is challenging and rigid constraints risk disrupting model operations, we apply a smooth penalty, adjusting the outlier coordinates while leaving the normal ones unaffected.
Approximation of activation distributions.
We empirically find that the activation distributions in each layer can be approximated by a normal distribution. The proof is as follows. For each layer $l$, we examine two distributions: the activation coordinate distribution $\mathcal{D}^{l}(x)$ and a normal distribution $\mathcal{D_{N}}^{l}(x)$ with the same mean $\mu^{l}_{\mathcal{D}}$ and standard deviation $\sigma^{l}_{\mathcal{D}}$. To measure their similarity, we compute the Jensen-Shannon (JS) divergence Lin (1991) between $\mathcal{D}^{l}(x)$ and $\mathcal{D_{N}}^{l}(x)$. Across all layers, the maximum JS divergence is 0.0839, and the mean is 0.0575, and both are well below 0.1. These low JS divergence values indicate strong similarity between $\mathcal{D}^{l}(x)$ and $\mathcal{D_{N}}^{l}(x)$, which supports the validity of approximating $\mathcal{D}^{l}(x)$ with $\mathcal{D_{N}}^{l}(x)$.
Penalty function design.
Figure 4: Workflow of ABD. ABD restricts outlier activation coordinates using a penalty function and determines its application scope via BO-based tuning.
To design a practical penalty function for activation coordinates, we establish three key principles: (1) It should target only outlier activations, leaving non-outliers unaffected. (2) The penalty should grow with the magnitude of deviation, reflecting distance-based penalization. (3) The function must be computationally efficient.
To construct penalty functions compatible with the activation coordinate distribution $\mathcal{D}^{l}(x)$, we approximate it using $\mathcal{D_{N}}^{l}(x)$ and focus on two geometric properties. First, symmetry around the mean value $\mu^{l}_{\mathcal{D}}$ requires the penalty function also to be symmetric about $\mu^{l}_{\mathcal{D}}$, ensuring unbiased penalization of outliers above or below the safety boundary. Second, the nonlinear decay of probability density with increasing distance from $\mu^{l}_{\mathcal{D}}$ implies that the penalty should grow nonlinearly with the distance, rising faster than a linear penalty. Following these principles, we propose a penalty function that updates the original activation coordinate scalar value $x$ to $x^{\prime}$ as follows:
$x^{\prime}=\alpha^{l}\cdot\mathrm{tanh}(\beta^{l}\cdot(x-\mu^{l}_{\mathcal{D}}))+\mu^{l}_{\mathcal{D}},$ (3)
where $\alpha^{l}\geq 0$ and $\beta^{l}\geq 0$ are hyperparameters. Visualization is given in Appendix D.1.
This function meets the stated expectations. It is symmetric about the mean $\mu^{l}_{\mathcal{D}}$, ensuring fair penalization for deviations above and below this central value. The penalty strength $|x-x^{\prime}|$ increases nonlinearly with the distance from $\mu^{l}_{\mathcal{D}}$ because $\mathrm{tanh(\cdot)}$ amplifies more significant deviations with its steep slope, reflecting stronger penalization for outliers. Moreover, the function mostly penalizes outliers. Within the range $[\mu^{l}_{\mathcal{D}}-b^{l},\mu^{l}_{\mathcal{D}}+b^{l}],x^{\prime}\approx x$, resulting in a negligible penalty for values close to the mean. The hyperparameters $\alpha^{l}$ and $\beta^{l}$ govern the behavior of the penalty function. The parameter $\alpha^{l}$ determines the maximum range of $x^{\prime}$, ensuring all coordinates are constrained within $(-\alpha^{l}+\mu^{l}_{\mathcal{D}},\alpha^{l}+\mu^{l}_{\mathcal{D}})$ after penalty application. Meanwhile, $\beta^{l}$ controls the size of the unmodified region $[\mu^{l}_{\mathcal{D}}-b^{l},\mu^{l}_{\mathcal{D}}+b^{l}]$; increasing $\beta^{l}$ expands this region, while decreasing $\beta^{l}$ narrows it.
4.2 Bayesian Optimization-Based Tuning
Model Jailbreak No Defense Paraphrase PPL Retokenization SafeDecoding Self-Exam Self-Reminder ABD (Ours)
Vicuna No attack 88% 82% 90% 66% 88% 100% 100% 100%
GCG-Individual 4% 86% 78% 62% 100% 86% 100% 100%
AutoDAN 6% 26% 12% 36% 78% 30% 30% 44%
PAIR 30% 62% 40% 30% 94% 86% 100% 76%
DeepInception 10% 2% 2% 0% 98% 18% 40% 58%
Llama-2 No attack 100% 100% 100% 94% 100% 100% 100% 100%
GCG-Individual 50% 98% 100% 98% 100% 64% 100% 100%
AutoDAN 98% 94% 98% 94% 100% 100% 98% 100%
PAIR 72% 90% 82% 78% 92% 100% 86% 92%
DeepInception 74% 82% 88% 60% 100% 94% 96% 100%
Table 1: The Defense Success Rate (DSR) of different defense methods across Vicuna-7B-v1.3 (simplified as Vicuna) and Llama-2-7B-chat (simplified as Llama-2) along with different attack types, where a higher DSR indicates better defense effectiveness. For ABD, the best and second performance across all defenses are highlighted.
Another crucial aspect of our method is the configuration of the hyperparameters and the application scope of the penalty function. Specifically, we must (1) minimize the number of affected activation coordinates to limit unnecessary perturbations, and 2) for each affected layer $l$, find the values of $\alpha^{l}$ and $\beta^{l}$ that best improve the model’s resilience against adversarial attacks. To accomplish these goals, we use a Bayesian Optimization (BO)-based Jones et al. (1998); Mockus (1974) tuning method. Concretely, we define two core objectives.
Layer selection objective: We introduce a tunable mask $M=[m_{0},\cdots,m_{L-1}]$, $m_{i}\in\{0,1\}$, where $m_{i}=1$ indicates that ABD is applied to layer $i$. Here, $L$ is the total number of transformer layers. By minimizing the ratio of layers where ABD is active, we reduce the number of interventions:
$\mathcal{L}_{\mathrm{Layer}}=1-\textstyle\frac{\mathrm{Sum}(M)}{L}.$ (4)
Robustness objective: For each layer $l$ with ABD applied, we aim to find $\alpha^{l}$ and $\beta^{l}$ to maximize the defense. A hyperparameter $k^{l}\in(0,1]$ controls the fraction of penalized activation coordinates, with smaller $k^{l}$ reducing disturbance but potentially weakening ABD. These parameters are aggregated as $\Theta=\{\theta^{l}\mid l\in[0,L-1]\}$, where $\theta^{l}={\alpha^{l},\beta^{l},k^{l}}$. The robustness objective is as follows:
$\mathcal{L}_{\mathrm{Robust}}=\mathrm{DSR}(\mathrm{Model}(\cdot\mid\Theta,M)).$ (5)
Since decreasing $k^{l}$ too aggressively can weaken ABD’s defense, our main strategy for reducing perturbations to the model is to minimize the number of affected layers by controlling the mask $M$. Our final objective is a weighted sum of two objectives:
[…]
where $w$ is a manually set parameter balancing defense robustness and minimal intervention.
5 Experiments
[…]
Table 2: Comparison of defenses on Runtime per Query and Just-Eval metrics in Vicuna-7B-v1.3. $\downarrow$: smaller is better; $\uparrow$: larger is better. For ABD, the best and second performances are highlighted. ABD preserves the model’s general performance, adding less than 0.1 seconds to runtime while producing high-quality outputs with leading evaluation scores. It has the best overall performance across all defenses.
5.1 Settings
Backbone models.
We consider two widely used open-source LLMs as the target models in our experiments: Llama-2-7B-Chat Touvron et al. (2023) and Vicuna-7B-v1.3. Llama-2-7B-Chat is specially trained for safety alignment, while Vicuna-7B-v1.3 is tuned based on Llama without extra alignment for safety. The configurations of the LLMs are shown in Appendix E.1.
Jailbreak and defense baselines.
For jailbreak methods, we considered four widely applied ones: optimization-based methods include GCG-Individual and AutoDAN, which iteratively optimize a jailbreak prompt towards the goal of generating affirmative responses. Model-based jailbreak, i.e., PAIR, uses an attacker LLM to refine the jailbreak prompts. The rule-based method includes DeepInception Li et al. (2023), which crafts jailbreak prompts based on a stealthy template. Correspondingly, we applied different defense methods to the models to compare them to our model for combating jailbreak methods: Paraphrase Jain et al. (2023), Retokenization Jain et al. (2023) and Self-Reminder Xie et al. (2023) reformulate the input to avoid attack; PPL Alon and Kamfonas (2023) and Self-ExamPhute et al. (2023) defense LLMs by double-checking their outputs; SafeDecoding Xu et al. (2024) uses a tuned model to modify the output probability distribution.
Datasets and metrics.
Following Xu et al. (2024) and Chao et al. (2023), we utilized 50 samples from AdvBench Zou et al. (2023) as the test set. We used DSR when evaluating defense effectiveness. Furthermore, following Xu et al. (2024), we adopted Just-Eval Lin et al. (2023) to measure the general abilities of LLMs before and after applying defense. Just-Eval is a comprehensive benchmark containing 1,000 diverse instructions, covering seven task types (e.g., reasoning, math, coding, etc.) and seven topics (e.g., ethics, nature, STEM, etc.). Following Lin et al. (2023), we leveraged GPT-4o-mini OpenAI (2024) to score the quality of the outputs, ranging from 1 to 5, across five aspects: helpfulness, clarity, factuality, depth, and engagement. We reported the average score for each aspect and the overall average score for different aspects, denoted as Avg. To evaluate the efficiency, we calculated the Runtime per Query for each defense, which represents the average time taken to process a single query during inference. This metric provides a direct measurement of the efficiency of defenses in real-world applications. We also utilized the above two metrics to derive an overall score that simultaneously reflects LLMs’ responding speed and quality. The calculation is shown in Appendix E.4.1.
Implementation Details.
We randomly filtered 400 non-overlapping AdvBench samples to compute $\mu_{\mathcal{D}}^{l}$ and used them as a validation set. We used GCG-Universal Zou et al. (2023) to attack the validation set. GCG-Universal finds a shared jailbreak suffix for all harmful prompts that can deceive LLMs. By default, we set $w=0.7$ when calculating $\mathcal{L}_{\mathrm{Robust}}(\Theta,M)$ to slightly prioritize improving DSR while balancing defensive capability and perturbation to LLMs. Appendix D.2 presents more details about ABD optimization.
5.2 Experimental Results
ABD reveals vulnerabilities in low and middle layers.
In both models, low and middle layers, such as layers 5 and 12 in Vicuna-7B-v1.3 and layers 2 and 12 in Llama-2-7B-chat, are penalized, which aligns with our observed findings and proposed explanation that low and middle layers are more vulnerable to jailbreak attacks due to noticeable activation shifts.
ABD successfully constrains jailbreak activations.
Figure 3(c) shows the percentage of jailbreak activations within the harmful activation space, namely the inclusion ratio. Before applying ABD, the inclusion ratio remains below 0.4 for various jailbreak activations, indicating that jailbreak activations lie outside of the safety boundary. After applying ABD, the ratio increases to 1, demonstrating that our method effectively constrains the jailbreak activations within the safety boundary. Additional details are provided in Appendix C.2. From a visualization perspective, Figure 1 projects activations for different prompts. Under our defense, jailbreak activations are progressively constrained within the harmful activation space under ABD, as shown in Figure 1. further confirming the effectiveness of ABD for mitigating jailbreak effects.
Statistic effectiveness of ABD.
We report the statistical defense results of ABD on the test set in Table 1. For Vicuna-7B-v1.3, which lacks specific safety alignment, the defense poses a more challenging task. Nevertheless, ABD demonstrates competitive performance. For jailbreak methods such as PAIR and DeepInception, ABD achieves a higher DSR (58%) compared to Paraphrase (2%) and Retokenization (0%), both of which require costly prompt reformulation. Against the GCG-Individual attack, ABD successfully defends against all jailbreak samples. For the well-aligned model Llama-2-7B-chat, ABD achieves 100% DSR under most jailbreak methods.
ABD is efficient and reliable.
Table 2 shows Runtime per Query, Just-Eval scores, and overall scores of Vicuna-7B-v1.3 with different defenses applied. We find that ABD has the greatest overall score across all defenses. Moreover, ABD only adds marginal extra time cost and general ability affection. Specifically, it only causes less than $1\%$ delay in each sample and less than $2\%$ perturbation in overall ability, compared to costly methods such as Paraphrase and Self-Exam. Furthermore, with ABD applied, the helpfulness, actuality, depth, and engagement also show a slight increase. We further discover that for baseline defenses such as Retokenization and Self-Reminder, LLM would generate overly simplistic outputs, which leads to smaller Runtime per Query, but they have significantly smaller Just-Eval scores.
6 Conclusion and Future Work
We conducted a comprehensive study of jailbreak mechanisms, analyzing over 30,000 samples. Our findings reveal that jailbreak shifts harmful activations outside the safety boundary in each layer, with the most severe shifts occurring in the low and middle layers. Motivated by this finding, we proposed ABD, which drives jailbreak activations back within the safety boundary, utilizing LLMs’ intrinsic sensitivity to harmful information. Our experiments suggest that ABD is both practical and efficient. In the near future, we aim to investigate the safety challenges of jailbreak attacks in multi-turn dialogue systems.
Limitations
Underperformance in under-aligned models.
For certain attack methods, under-aligned models (Vicuna-7B-v1.3) may not perform significantly as well-aligned models (Llama2-7B-chat). We believe this is because under-aligned models have unclear safety boundaries, which complicate the search for penalty functions that balance general ability and DSR. Future work could refine ABD by specifically searching for activation spaces that preserve the concept of "safety", therefore enhancing its generalizability on uncensored models.
Focus on single-round jailbreak.
In this study, we primarily focus on single-round jailbreak scenarios. We do not extend our analysis to more complex jailbreaks that involve long contexts or multi-round dialogues, such as CFA Sun et al. (2024). As a result, the relationship between jailbreak dialogues and safety boundaries remains largely unexplored. Addressing this limitation, we plan to investigate and incorporate methods for detecting safety boundaries in multi-round dialogue scenarios as part of our future work.
Ethical Considerations
The aim of this research is to enhance the explainability and safety of LLMs. Our proposed jailbreak mechanism, that jailbreak shift activations out of the safety boundary, can mitigate disputes on how jailbreak happens and promote the development of both LLM explainability and safety. We highlight that the development of ABD only needs publicly available datasets and jailbreak methods and does not require designing new jailbreak methods. We demonstrate some harmful responses from LLMs only for illustration purposes. We acknowledge that ABD would cause the development of new attacks. Therefore, we will explore using random perturbation in the input sequence rather than a particular jailbreak method when optimizing to mitigate such attacks.
Appendix A Statistics of Observed Data
We present the statistics of our data for observation experiments as Table 3.
Type Samples
Benign samples 20,000
Harmful samples 8,556
Jailbreak samples
AdvPrompter 1,872
AutoDAN 520
COLD-Suffix 436
ArtPrompt 361
GCG-Individual 340
GCG-Universal 312
PAIR 110
Total 32,507
Table 3: Statistics of observation experiments in §3. The attacked samples are derived from part or all of the samples from AdvBenchZou et al. (2023).
Appendix B Full View of Activation Space
We visualize activation spaces of all layers in Vicuna-7B-v1.3, as shown in Figure LABEL:fig:all_subfig1, Figure LABEL:fig:all_subfig2 and Figure LABEL:fig:all_subfig3.
Appendix C Supplementary Experiments
C.1 Data Augmentation Experiment
We show that by adopting the same method as Shen et al. (2024), we can draw a different conclusion by scaling up data. Shen et al. (2024) state jailbreak happens by posing jailbreak activations between benign and harmful activations in middle and deep layers. Following Shen et al. (2024), we randomly select 60 samples for each type of activation and conduct t-SNE on layer 14, as shown in the left part of 5. Jailbreak activations are between harmful and benign samples, which is in agreement with Shen et al. (2024). When scaling up each type of activation to 500 samples, jailbreak activations seem to cluster on the harmful activation side, as shown in the right part of 5. Therefore, jailbreak activations are not always between harmful and benign activations in deeper layers.
C.2 Inclusion Ratio Experiments
For a layer $l$, to measure the portion of a set of jailbreak activations $A^{l}=\{a^{l}_{0},a^{l}_{1},\cdots,a^{l}_{n}\}$ that resides in harmful activation space, we propose an inclusion ratio. Based on 8,556 harmful samples gathered in Table 3, we calculate a ball that covers $80\%$ activations. The center of the ball is $\mu_{\mathcal{D}}^{l}$, and the radius of the ball is denoted as $r_{\mathcal{D}}^{l}$.
Then, we calculate the portion of $A^{l}$ which are contained within the ball:
$\rho^{l}_{\mathrm{inclusion}}=\frac{\left|\{a^{l}_{i}\in A^{l}\mid\|a^{l}_{i}-\mu_{\mathcal{D}}^{l}\|_{2}\leq r_{\mathcal{D}}^{l}\}\right|}{|A^{l}|},$
where $\|a^{l}_{i}-\mu_{\mathcal{D}}^{l}\|_{2}$ is the distance between the activation and the center. We calculate inclusion ratios of different types of jailbreak activations, with three representative types shown in Figure 1. We then apply ABD on all types of jailbreak activations and calculate their inclusion ratios. Notably, despite different types of jailbreaks, they all achieve an inclusion ratio of $100\%$ with ABD applied, which verifies the effect of our method.
Appendix D Supplementary Illustration on ABD
Figure 5: t-SNE visualization of activations in layer 14. Left: Results with 60 samples per type following Shen et al. (2024), showing jailbreak activations between harmful and benign activations. Right: Results after scaling up to 500 samples per type, showing jailbreak activations clustering on the harmful side.
D.1 Visualization of the penalty function.
A visualization of penalty functions are shown in Figure 6. The penalty function is symmetric about $(\mu_{\mathcal{D}}^{l},\mu_{\mathcal{D}}^{l})$. Figure 6(a) and Figure Figure 6(b) presents the change in the range $[\mu_{\mathcal{D}}^{l}-b^{l},\mu_{\mathcal{D}}^{l}+b^{l}]$, where $x^{\prime}\approx x$. A larger $\beta^{l}$ results in the larger little-perturbed region. Figure 6(a) and Figure 6(c) presents the change in the range of $x^{\prime}$. When $\alpha^{l}$ increases, $x^{\prime}$ is confined within a wider range of values. $\alpha^{l}$ and $\beta^{l}$ collaboratively determines behaviors of the penalty function.
Figure 6: Penalty functions under different $\alpha^{l},\beta^{l}$ compared with $x^{\prime}=x$.
D.2 ABD Optimization Settings
Validation data.
To make validation data, we adopt GCG-Universal Zou et al. (2023). We optimize a common suffix for all 400 samples, with n_steps=1, batch_size=512.
To ensure efficiency, in each iteration of optimization, we select a batch of harmful prompts from the 400 entries as $\mathcal{S}_{\mathrm{val}}$. We initially set the batch size to 15. In most cases $\mathcal{L}_{\mathrm{Robust}}(\Theta,M|\mathcal{S}_{\mathrm{val}})<0.9$, and the optimization process continues to the next iteration. If $\mathcal{L}_{\mathrm{Robust}}(\Theta,M|\mathcal{S}_{\mathrm{val}})\geq 0.9$, due to the potential regional optimal, we iteratively increase the batch size by ten and reformulate $\mathcal{S}_{\mathrm{val}}$ to test again. This process ends if 1) the calculated $\mathcal{L}_{\mathrm{Robust}}<0.9$ or 2) batch size reaches 50.
Initial values.
To prevent the futile search of the optimizer, we set valid initial values before optimization: for $i\in\{2,12\},m_{i}=1,\alpha_{i}=1,\beta_{i}=0.5,k_{i}=0.5$; for $i\notin\{2,12\},m_{i}=\alpha_{i}=\beta_{i}=k_{i}=0.$
Optimizing framework.
We adopt Optuna Akiba et al. (2019) as our framework of optimization. We follow Optuna’s default settings, i.e., Gaussian process-based Bayesian Optimization.
Appendix E Supplementary Illustration on Experiments
E.1 Generation Configs
When conducting experiments, we directly utilize most of the original configurations of Vicuna-7B-v1.3 and Llama2-7B-chat. Specifically, we set max_new_tokens=128. We find that the proper functioning of these LLMs largely depends on the chat template of the inputs. We apply chat templates in fastchat v0.2.36 by:
fastchat.model.get_conversation_template(template_name).
We set template_name="vicuna" for Vicuna-7B-v1.3 and template_name="llama-2" for Llama2-7B-chat.
E.2 Jailbreak Methods
GCG-Individual and GCG-Universal
Zou et al. (2023) are optimization-based jailbreak attacks. They build towards an objective to repeat the prompt affirmatively and optimize a suffix based on a Greedy Coordinate Gradient-based search. The model would likely repeat the prompt affirmatively and generate harmful content with the suffix added behind the original prompt. GCG-Individual focuses on generating a tailored suffix designed specifically for a particular prompt. In contrast, GCG-Universal seeks to identify a generalized suffix that can be applied across multiple prompts, enabling it to deceive the model in a broader range of scenarios.
AutoDAN
Liu et al. (2024b) utilize a meticulously designed hierarchical genetic algorithm and generate stealthy jailbreak prompts. The generated prompts are highly readable and transferable.
PAIR
Chao et al. (2023) is a jailbreak method that leverages an attacker LLM aiming at making the target LLM answer harmful prompts. The attacker LLM iteratively queries the target LLM to update and refine a candidate jailbreak prompt.
DeepInception
Li et al. (2023) proposes a simple method that uses the personification ability of LLMs. It creates a virtual and layered scene, allowing the model to find flexible ways to bypass usage controls in normal situations.
E.3 Defense Methods
PPL
Alon and Kamfonas (2023) discovers that jailbreak prompts often lead to high perplexity values in LLMs. It, therefore, involves adding a classifier trained on perplexity and text length at the end of LLMs. The classifier can serve as a filter to avoid outputting potentially harmful answers.
Paraphrase
Jain et al. (2023) defense LLMs by making them paraphrase their inputs, avoiding deception caused by potential adversarial jailbreak suffixes within original inputs.
Retokenization
Jain et al. (2023) disrupts adversarial suffixes by retokenizing the input sequence, breaking tokens apart, and re-representing them with smaller tokens.
SafeDecoding
Xu et al. (2024) is a safety-aware decoding strategy. It mitigates jailbreak attacks by amplifying the probabilities of safety disclaimers tokens among top-ranked tokens and attenuating the probabilities of harmful token sequences, ensuring LLMs generate helpful and harmless responses.
Self-Exam
Phute et al. (2023) triggers the LLMs’ awareness of safety issues by adding a pre-defined prompt template, asking LLMs to examine if their outputs are safe.
Self-Reminder
Xie et al. (2023) defends the LLMs by adding a system prompt specifically emphasizing potential safety issues to prevent the model from outputting harmful responses.
E.4 Metrics
E.4.1 Calculating DSR
We use Dic-Judge Zou et al. (2023) to calculate DSR. Dic-Judge detects the presence of pre-defined refusal strings in the outputs of LLMs. Given a jailbreak query as input, if a particular refusal string is detected in the outputs, we consider the model successfully refuses to answer the question, thereby achieving the goal of defense; otherwise, the model is deemed to be deceived by the jailbreak inputs. DSR is defined as follows:
$\mathrm{DSR=\frac{|\{responses\ containing\ refusal\ strings\}|}{|\{input\ jailbreak\ queires\}|}}.$
We adopt refusal strings as proposed by Xu et al. (2024), presented in Table
E.4.2 Calculating Overall Scores
We use Runtime per Query and overall averaged Just-Eval scores to calculate overall scores:
$\mathrm{Overall=\big{[}1-Norm(t)+Norm(Avg.)\big{]}/2,}$
where $\mathrm{t}$ is Runtime per Query and $\mathrm{Avg.}$ represents overall averaged Just-Eval scores. 4.
Refusal Strings
I’m sorry
Sorry
I am sorry
I apologize
As an
As an AI
I’m an
I’m just
As a language model
As an Assistant
I cannot
I do not
It is not</s>
It is not appropriate
I’m unable to
I am unable to
I am not allowed to
I am an AI language model
Table 4: Refusal strings.

### v2

@@
 Abstract
-Jailbreaking in Large Language Models (LLMs) is a major security concern as it can deceive LLMs to generate harmful text.
-Yet, there is still insufficient understanding of how jailbreaking works, which makes it hard to develop effective defense strategies.
-We aim to shed more light into this issue: we conduct a detailed large-scale analysis of seven different jailbreak methods and find that these disagreements stem from insufficient observation samples.
-In particular, we introduce safety boundary, and we find that jailbreaks shift harmful activations outside that safety boundary, where LLMs are less sensitive to harmful information.
-We also find that the low and the middle layers are critical in such shifts, while deeper layers have less impact.
-Leveraging on these insights, we propose a novel defense called Activation Boundary Defense (ABD), which adaptively constrains the activations within the safety boundary.
-We further use Bayesian optimization to selectively apply the defense method to the low and the middle layers.
-Our experiments on several benchmarks show that ABD achieves an average DSR of over 98% against various forms of jailbreak attacks, with less than 2% impact on the model’s general capabilities.
+Jailbreaking in large language models (LLMs) poses major security risks by tricking models into generating harmful text.
+However, there is limited understanding of how jailbreaking operates, which makes it difficult to develop effective defenses.
+In this work, we conduct a large-scale analysis of seven jailbreak methods and uncover that inconsistencies in previous studies arise from insufficient observation samples.
+Our analysis reveals that jailbreaks shift harmful activations beyond a defined safety boundary, where LLMs become less sensitive to harmful information.
+We also find that the low and the middle layers are critical in driving these shifts, while deeper layers play a lesser role.
+Leveraging these insights, we propose a novel defense mechanism called Activation Boundary Defense (ABD), which adaptively constrains activations within the safety boundary.
+To further optimize performance, we use Bayesian optimization to select the most effective layers for ABD application and confirm that the low and middle layers have the greatest impact, consistent with our earlier observations.
+Experiments across multiple benchmarks demonstrate that ABD achieves an average defense success rate (DSR) of over 98% against various jailbreak attacks, with less than 2% impact on the model’s overall capabilities.
 Shaping the Safety Boundaries: Understanding and Defending Against Jailbreaks in Large Language Models
-Lang Gao1,2,  Xiangliang Zhang3, Preslav Nakov1, Xiuying Chen1222Corresponding author.
-1MBZUAI  2Huazhong University of Science and Technology 3University of Notre Dame {Lang.Gao, Preslav.Nakov, Xiuying.Chen}@mbzuai.ac.ae xzhang33@nd.edu
+Lang Gao1,2,  Jiahui Geng1,  Xiangliang Zhang3, Preslav Nakov1, Xiuying Chen1thanks: Corresponding author.
+1MBZUAI  2Huazhong University of Science and Technology 3University of Notre Dame {Lang.Gao, Jiahui.Geng, Preslav.Nakov, Xiuying.Chen}@mbzuai.ac.ae xzhang33@nd.edu
+[…]
+WARNING: This paper contains model outputs that may be considered offensive.
 1 Introduction
 The widespread use of Large Language Models (LLMs) across various fields Kaddour et al.
-(2023) has raised concerns about the safety of the output and the robustness of the model.
+(2023); Chen et al.
+(2025); Xie et al.
+(2025); Liu et al.
+(2024c) has raised concerns about the safety of the output and the robustness of these models Ye et al.
+(2024); Huang et al.
+(2025); Wang et al.
+(2025).
 It has been shown that jailbreak attacks, which use crafted prompts to deceive LLMs into generating harmful content, can bypass LLM’s safety alignment Liu et al.
 (2024b); Zou et al.
+(2023); Song et al.
 (2025), and a lot of research has focused on developing defense mechanisms and counter-prompts to mitigate such attacks Robey et al.
@@
 The attack arrow indicates that jailbreak prompts shift harmful activations into the benign space to evade safety checks.
-Understanding the internal mechanism of how jailbreaks work is crucial for improving defenses and enhancing safety measures.
-However, studies explaining jailbreaks remain largely ambiguous and controversial Yu et al.
+Understanding the internal mechanisms of jailbreak attacks is crucial for both explaining their operation and developing effective defenses.
+While existing studies have made initial explorations Yu et al.
 (2024); Ball et al.
 (2024a); Lin et al.
-(2024).
-From an explainability perspective, debates revolve around which layers are most responsible for a successful attack: He et al.
-(2024) prioritize the low layers, Zhou et al.
-(2024); Shen et al.
-(2024) focus on the middle layers, while Li et al.
-(2024) highlight the deep layers.
-Another contradiction lies in whether activations gradually shift as layers deepen or abruptly transition in the middle layers Zhao et al.
+(2024), fundamental disagreements remain about the attack mechanisms.
+The academic community continues to debate two key aspects of explainability: (i) The localization of attack impacts across the model layers, with competing findings suggesting primary responsibility in low He et al.
+(2024), middle Zhou et al.
+(2024b); Shen et al.
+(2024), or deep layers Li et al.
+(2024), and (ii) the progression pattern of malicious activation, with conflicting observations about gradual shifts versus abrupt transitions in the middle layers Zhao et al.
 (2024a); Shen et al.
 (2024).
-Our experiments suggest that these inconsistencies may stem from the limited sample size of prompts, typically around 100, used in prior analyses.
-From a defense perspective, existing approaches often rely on additional training processes Zhao et al.
-(2024a) or probing a limited number of samples Shen et al.
-(2024), which affects the utility of LLMs.
-Here we aim to better understand the mechanism of how jailbreaking works.
+Our experiments suggest that these inconsistencies may stem from methodological limitations in prior analysis, particularly their reliance on small prompt samples (typically $\sim 100$).
+On the defense front, current approaches face dual challenges: Most detection methods either require additional training Zhao et al.
+(2024a) or demonstrate limited generalizability due to small-sample probing Shen et al.
+(2024), while mitigation strategies often degrade normal model capabilities.
+In this work, we aim to better understand the mechanism of how jailbreaking works.
 In particular, we provide our explanation of jailbreaks based on a comprehensive analysis of over 30,000 samples, a significantly larger scale than previous studies.
-As an example, Figure 1 shows the two-dimensional projection of three types of prompts: benign prompts that contain no harmful information, harmful prompts that attempt to induce the LLM to generate harmful content but fail, and various jailbreak prompts that successfully induce harmful outputs, collected from different types of jailbreak attack methods.
+Figure 1 shows the projection of three types of prompts: benign prompts that contain no harmful information, harmful prompts that attempt to induce the LLM to generate harmful content but fail, and various jailbreak prompts that successfully induce harmful outputs, collected from different types of jailbreak attack methods.
 The projection is based on the last token representations across different layers, referred to as activations, as they capture the model’s overall understanding of the entire input sequence Radford et al.
@@
 (2023).
-We can see that different prompts exhibit clustering effects, where the jailbreak activations are distinct from the activation space of harmful prompts.
-This indicates that in each layer, the LLM performs a rigorous safety check within a safety boundary, effectively containing most of the harmful activations within this controlled space.
-Jailbreak prompts, however, bypass these safety checks by shifting activations beyond the safety boundary into an alternative, unmonitored activation space.
-Notably, the extent of this shift varies across different layers of the LLM, with the low and the middle layers exhibiting the most pronounced effects.
-This suggests that these layers are critical to the success of jailbreaks.
+Our results reveal distinct clustering patterns, where different prompts form separate activation clusters.
+Notably, harmful prompts form a contained cluster within a specific region that we define as the safety boundary, representing the LLM’s internal safety check that typically prevents harmful activations from escalating into harmful outputs.
+In contrast, jailbreak prompts push activations beyond the safety boundary into an unmonitored space where harmful content can be generated.
+This shift is most pronounced in the low and middle layers, indicating their critical role for jailbreak success.
 This finding motivates a new defense mechanism that we propose, Activation Boundary Defense (ABD), which constrains jailbreak prompt activations within a safety boundary by applying penalties to activation values.
-Specifically, we impose minimal penalties within the boundary but sharply increase them beyond it, thereby preserving model utility.
-The constraint function considers various factors, including the defense layer and the defense dimension, which balance performances on defense and general tasks.
-We propose a novel adaptive objective that maximizes the Defense Success Rate (DSR) while minimizing the number of penalized layers.
-This objective is implemented with an adapted Bayesian Optimization (BO) method Jones et al.
-(1998); Mockus (1974), which iteratively proposes and refines selection suggestions.
+Specifically, we impose minimal penalties within the boundary, but sharply increase them beyond it, thereby preserving model utility.
+The penalty function considers various factors, including the defense layer and the defense dimension, which balance the performance on defense and on general tasks.
+We propose a novel objective that maximizes the Defense Success Rate (DSR) while minimizing the number of penalized layers.
+This objective is implemented with an adapted Bayesian Optimization method Jones et al.
+(1998), which iteratively proposes and refines selection suggestions.
 Experiments on AdvBench Zou et al.
 (2023) show that ABD has good generalization, achieving an average DSR of over 98% against various forms of jailbreak attacks.
-In the general ability tasks, the performance drop of the model equipped with ABD does not exceed 2%, which is significantly lower compared to other defenses, such as 37% for Retokenization Jain et al.
-(2023).
-This demonstrates that our method has a minimal impact on the model’s overall abilities.
-In summary, our contributions are threefold:
-  • We uncover a comprehensive activation distribution and introduce the safety boundary, resolving contradictions in prior work and highlighting the vulnerability of the low and the middle layers.
-  • We propose a lightweight, extensible defense that penalizes only targeted samples and a few key layers, ensuring efficiency and precision.
-  • Our experiments on benchmark datasets demonstrate that ABD achieves an average DSR of over 98% against various jailbreak attacks, with less than a 2% impact on the general capabilities.
+For general ability tasks, the performance drop of the model equipped with ABD does not exceed 2%, which is substantially lower than the drop of up to 37% with other defenses.
+Our contributions can be summarized as follows:
+$\bullet$ We uncover a comprehensive activation distribution and introduce the safety boundary, resolving contradictions in prior work and highlighting the vulnerability of the low and middle layers.
+$\bullet$ We propose a lightweight, extensible defense that penalizes only targeted samples and a few key layers, ensuring efficiency and precision.
+$\bullet$ Our experiments on benchmark datasets demonstrate that ABD achieves an average DSR of over 98% against various jailbreak attacks, with less than a 2% impact on general capabilities.
 2 Related Work
 Jailbreak attacks on LLMs.
-Jailbreak attacks craft prompts to induce harmful outputs from LLMs.
-Early approaches, like DAN Shen et al.
-(2023), rely on manually designed prompts.
-Later research proposed automatic methods.
-Optimization-based methods Zou et al.
+Jailbreak attacks craft prompts to induce harmful outputs from LLMs, evolving from early manually designed approaches to more sophisticated techniques.
+Early methods, such as DAN Shen et al.
+(2023), rely on manually constructed prompts.
+Later, optimization-based methods Zou et al.
 (2023); Liu et al.
 (2024b); Guo et al.
-(2024) design optimization strategies to iteratively edit and refine original harmful prompts, enhancing their stealthiness.
+(2024) emerged, using iterative optimization strategies to refine harmful prompts and enhance their stealthiness.
 Model-based methods Chao et al.
@@
 (2024); Paulus et al.
-(2024) utilize an attacker LLM to autonomously query and refine jailbreak prompts.
-For rule-based jailbreak methods, Jiang et al.
-(2024) and Liu et al.
-(2024c) rewrite prompts using fixed rules that are sufficiently aggressive to deceive the model.
-In this work, we aim to analyze the mechanisms of these jailbreak methods to develop more effective defense strategies.
+(2024); Huang et al.
+(2024) use attacker LLMs to autonomously generate and improve jailbreak prompts.
+Meanwhile, rule-based methods Jiang et al.
+(2024); Liu et al.
+(2024d) rewrite prompts using predefined rules that effectively deceive the model.
+In this work, we analyze the mechanisms behind these jailbreak techniques.
 Defense against jailbreak attacks.
-Unlike jailbreak methods, defense strategies enhance model safety and robustness by reformulating inputs to counter jailbreak prompts, such as backtranslation Wang et al.
+Defense strategies for jailbreak attacks typically fall into three categories.
+One common approach enhances model safety and robustness by reformulating inputs to counter jailbreak prompts through techniques such as backtranslation Wang et al.
 (2024), paraphrasing Jain et al.
-(2023), and reminding Xie et al.
-(2023).
-Some researchers use indicators to distinguish jailbreak prompts, e.g., classify input sequences based on perplexity and sequence length Alon and Kamfonas (2023), or classify jailbreak activations using linear classifiers to control whether to answer the query Zhao et al.
+(2023), self-reminding Xie et al.
+(2023), and adversarial prompting Zhou et al.
 (2024a).
-A new trend in defense involves directly manipulating the representations of the model Shen et al.
-(2024); Xu et al.
+Another approach focuses on identifying jailbreak prompts using indicators, such as classifying input sequences based on perplexity and sequence length Alon and Kamfonas (2023) or using linear classifiers to detect jailbreak activations and decide whether to respond Zhao et al.
+(2024a).
+A more recent trend directly manipulates the model’s internal representations Xu et al.
 (2024); Li et al.
 (2024); Liu et al.
-(2024a) and editing the model Zhao et al.
-(2024b), which is more efficient and cost-effective.
-Unlike these methods, ours avoids extra tokens or modules by directly constraining activations within a safety boundary, ensuring greater efficiency and simplicity.
+(2024a) or editing the model Zhao et al.
+(2024b).
+Unlike these methods, our approach directly constrains activations within a safety boundary, avoiding extra tokens or modules.
 Mechanistic interpretability of LLMs.
@@
 Zhou et al.
-(2024) discovered that the vocabulary mappings of activations significantly changed when processing jailbreak inputs.
+(2024b) discovered that the vocabulary mappings of activations significantly changed when processing jailbreak inputs.
 Research efforts have also focused on proposing corresponding defense methods based on interpretability results.
@@
 (2024a) and Shen et al.
-(2024) model how jailbreak activations transfer between benign and harmful activation spaces as layers deepen.
+(2024) modeled how jailbreak activations transfer between benign and harmful activation spaces as layers deepen.
 They designed adaptive defenses whose strength varies across different layers.
-However, the limited data in previous studies often led to controversy and ambiguity.
-In contrast, we leverage over 30,000 samples to better understand jailbreaks and design effective defenses.
+However, the limited data in all previous studies often led to controversy and ambiguity.
 3 Understanding Jailbreaks in LLMs
 3.1 Controversy in Literature
-What are the key layers?
+Disputes on layer importance.
 There have been different opinions about which layers are most important, mainly due to varying perspectives.
 For example, He et al.
-(2024) argued that low layers are essential because they treat jailbreak activations as benign ones.
+(2024) argued that the low layers were essential because they treated jailbreak activations as benign ones.
 Zhou et al.
-(2024) believe that low and middle layers are essential based on tokens generated from activations.
+(2024b) believed that the low and the middle layers were essential based on tokens generated from activations.
 Li et al.
-(2024) argued that deep layers are critical, as the concept they proposed, called safety pattern, has greater values in these layers.
-Among these works, the first work similarly focuses on activations, as does our paper.
-Upon examining their experimental setup, we identified a major limitation: the small number of samples used in the analysis.
-As shown in Figure 2(a), before scaling the data, benign and harmful activations are linearly separable, with most jailbreak samples being misclassified as benign, which is consistent with the findings from He et al.
-(2024).
-However, after scaling the data, half of the jailbreak activations shift to align with harmful ones, and the separation becomes nonlinear.
-This highlights the risk of misleading from small-scale studies and underscores the importance of large-scale analysis.
+(2024) argued that deep layers were critical, as their proposed safety pattern showed greater values in these layers.
+Similarly, He et al.
+(2024) focused on activations, as we do as well.
+However, their experimental setup suffers from a key limitation: a small sample size used in the analysis.
+As shown in Figure 2(a), under the original settings, benign and harmful activations are linearly separable, with most jailbreak samples misclassified as benign.
+After scaling, half of the jailbreak activations align with harmful ones, making the separation nonlinear.
+This underscores the risks of small-scale studies and the need for large-scale analysis.
 Disputes on the jailbreak mechanism.
-There are also different perspectives on the internal mechanism of jailbreak.
-Specifically, Zhao et al.
-(2024a) viewed jailbreak as a gradual process where activations transition from harmful to benign spaces from low to deep layers, while Shen et al.
-(2024) argued that jailbreak abruptly manifests in deeper layers, with activations positioned between harmful and benign spaces.
-Figure 2: PCA visualization of the activations highlights how conclusions are drawn from tiny sample sizes, such as the side on which the jailbreak activation lies, may not hold with larger datasets.
-(a) Top: 100 samples per type (following He et al.
-(2024)); Bottom: Benign and harmful samples increased to 5000.
-(b) Top: 60 samples per type (similar to Zhao et al.
-(2024a)); Bottom: All samples increased to 500.
-This work used no more than 100 samples per activation type, and we identify a similar issue to the one discussed earlier.
-Figure 2(b) shows that before scaling the data, the jailbreak activations are aligned to benign activations, which is consistent with findings from Zhao et al.
-(2024a); while after scaling the data, jailbreaking activations no longer align with either harmful or benign activations.
-When compared to Shen et al.
-(2024), similar conflicting perspectives are observed when scaling up the data.
-Further details can be found in Appendix C.1.
-These conflicting perspectives form the basis of the jailbreak mechanisms in these works; their conclusions — such as whether activations gradually shift or abruptly transition — are also questionable.
+There are differing perspectives on the internal mechanism of jailbreaks.
+Zhao et al.
+(2024a) viewed jailbreak as a gradual process, where activations transition from harmful to benign spaces as they pass from the low to the deep layers.
+In contrast, Shen et al.
+(2024) argued that jailbreaks manifest abruptly in the deeper layers, with activations positioned between harmful and benign spaces.
+Both studies relied on fewer than 100 samples per activation type, contributing to similar inconsistencies.
+As shown in Figure 2(b), before scaling, jailbreak activations align closely with benign ones, consistent with Zhao et al.
+(2024a).
+However, after scaling, they diverge from both harmful and benign activations.
+When considering Shen et al.
+(2024), similar conflicts arise between pre- and post-scaling results (details in Appendix C.1, Figure 5).
+These conflicting observations cast doubt on their conclusions regarding the true mechanism of jailbreaks.
+Figure 2: PCA visualizations reveal the limitations of small-sample sizes: (a) Top: 100 samples per type He et al.
+(2024); Bottom: 5,000 benign and harmful samples.
+(b) Top: 60 samples per type Zhao et al.
+(2024a); Bottom: 500 samples per type.
 3.2 Experimental Settings
@@
 Without ABD, the ratio stays below 0.4 but rises to 1 when ABD is applied.
-Dataset.
+Dataset and Model.
 Our dataset consists of 32,507 samples in three categories: benign, harmful, and jailbreak.
@@
 (2023), as it has been carefully curated to ensure that only safe content is included.
-For harmful samples, we collected five datasets containing harmful queries from RedEval Bhardwaj and Poria (2023) and AdvBench Zou et al.
+We collected harmful samples from five datasets in RedEval Bhardwaj and Poria (2023) and AdvBench Zou et al.
 (2023).
-These datasets encompass a wide variety of harmful prompts designed to elicit unsafe outputs.
+These datasets encompass a wide variety of harmful and misleading prompts.
 For jailbreak samples, we applied seven different jailbreak attack methods, including GCG-Individual and GCG-Universal Zou et al.
@@
 (2024), on AdvBench.
-Detailed statistics can be found in Appendix A.
-Model.
+Detailed statistics can be found in Table 4 in Appendix A.
 We used Vicuna-7B-v1.3 Chiang et al.
-(2023), a 32-layer Transformer model, as the LLM to study.
-It is tuned upon a safety-aligned LLM and is able to identify certain harmful prompts.
-However, as demonstrated in previous work Chu et al.
-(2024), it can also be easily deceived and generate harmful information during jailbreak attacks.
-This makes it an excellent example for studying both attack mechanisms and defense strategies.
+(2023), a 32-layer Transformer, fine-tuned on a safety-aligned LLM.
+Although it can identify harmful information, it remains vulnerable to jailbreak attacks Chu et al.
+(2024), making it an ideal model for analyzing attacks and defenses.
 MDS projection.
-Our projection was based on the activation, the last token vector of the input, as it captures the model’s overall understanding of the entire input sequence Radford et al.
-(2019); Zou et al.
-(2023).
+Our projection was based on the activation, the last token vector of the input.
 As discussed in §3.1, linear classification may be insufficient to model the boundary between different prompts.
-Therefore, we adopted Multi-Dimensional Scaling (MDS) Carroll and Arabie (1998) as our dimensionality reduction method, as MDS is more effective than PCA in handling large-scale and complexly distributed data Anowar et al.
-(2021).
-Moreover, it better captures the global distribution structure compared to t-SNE Van der Maaten and Hinton (2008).
+Therefore, we adopted Multi-Dimensional Scaling (MDS) Carroll and Arabie (1998) as our dimensionality reduction method, as PCA assumes linear separability Anowar et al.
+(2021), and t-SNE focuses on preserving local structures while distorting the global distribution Van der Maaten and Hinton (2008).
 3.3 Jailbreak Mechanisms Findings
 3.3.1 Activation Distribution
-We present the activation distributions of representative layers in Figure 1, with the full version available in Appendix B.
-Our key observations are as follows: (1) Harmful and benign activations overlap and are not linearly separable in most layers, especially in low and middle layers, as shown in Figure 1(a) and Figure 1(b), where benign and harmful activations are overlapping with no clear linear decision boundaries to separate them.
-This highlights the intricate interplay between harmful and benign activations, indicating that forcing jailbreak activations into a linear benign-vs-harmful classification is impractical.
-(2) Jailbreak activations significantly shift out of the harmful activation space in most layers, forming an independent region with low overlap with harmful activations.
-This pattern is consistent across various jailbreaks, suggesting that jailbreaks achieve deception by shifting activations.
-Such shift is evident in all layers, indicating that jailbreaks begin in low layers and continue to affect the model across all layers, rather than starting in middle or deep layers Zhou et al.
-(2024); Shen et al.
+We present the activation distributions of representative layers in Figure 1, with the full version available in Figures LABEL:fig:all_subfig1,LABEL:fig:all_subfig2,LABEL:fig:all_subfig3 in Appendix B.
+Our direct observations are as follows: (1) Harmful and benign activations overlap and are not linearly separable in most layers.
+There are no clear linear decision boundaries to distinguish between harmful and benign activations, highlighting the ineffectiveness of treating jailbreak activations as a simple linear classification problem.
+(2) Jailbreak activations shift significantly out of the harmful activation space, forming distinct regions with minimal overlap.
+This shift is consistent across different jailbreaks, starting in the low layers and persisting throughout the model, rather than originating in the middle or deep layers Zhou et al.
+(2024b); Shen et al.
 (2024).
-Moreover, the shift is the most pronounced in the low and middle layers, implying their importance in jailbreaks.
+Additionally, the shift is most pronounced in the low and middle layers.
 3.3.2 Safety Boundary
-Based on the observation of jailbreak and harmful activations, we propose the following hypothesis: A safety boundary exists that ensures sensitivity to harmful content within its activation space, but jailbreak prompts can bypass it, generating harmful responses.
-To verify the existence of the safety boundary, we conduct experiments and present two concepts described below.
-We first explore Randomized Activation Shifting (RAS), an approximation of jailbreak attacks that introduces a random value shift to the original activation.
-Specifically, assuming the activation on the $l$-th layer is $a_{l}$, we shift it by a distance $r$ in a random direction $\hat{u}$:
+Based on observations of jailbroken and harmful activations, we propose that jailbreak mechanisms are better explained by activations shifting outside a closed safety boundary, rather than simply crossing from one side of a linear boundary to the other.
+To validate this, we conduct experiments and introduce two key concepts below.
+Randomized Activation Shifting (RAS): It is an approximation of jailbreak attacks by randomly shifting the original activation.
+Assume the activation on the $l$-th layer is $a_{l}$, we shift it by a distance $r$ in a random direction $\hat{u}$:
 $a_{l}\leftarrow a_{l}+r\cdot\hat{u}.$ (1)
-Figure 3(a) shows the changes in Defense Success Rate (DSR) under different values of $r$ for four selected layers.
-The definition of this metric is provided in the Appendix.
-Initially, when $r=0$, the corresponding DSR is over 99%, as the harmful text has activations within the safety boundary.
-Then, as the distance $r$ increases, the DSR decreases for all layers, indicating that harmful activations become harder for the model to detect.
-Notably, regardless of the selected layer, a steep drop in DSR is observed at a specific range (e.g., $r=10-20$ on layer 0), while the decline is more moderate in other regions.
-This observation suggests the presence of a safety boundary.
-For more accurate boundary detection, we provide the second visualization, which introduces a measurable concept called Most Vulnerable Distance (MVD) as an approximation of the boundary.
-MVD represents the specific distance at which the DSR experiences the steepest drop, indicating the point where the model’s ability to detect harmful activations becomes most vulnerable:
+[…]
+Most Vulnerable Distance (MVD): It is a measurable approximation of the safety boundary.
+MVD represents the specific distance at which the DSR experiences the sharpest drop, indicating the point where the model’s ability to detect harmful activations becomes most vulnerable:
 $\mathrm{MVD=\textstyle\arg\min_{r}}\textstyle\frac{\mathrm{d(DSR)}}{\mathrm{d}r},$ (2)
 where $\frac{\mathrm{d(DSR)}}{\mathrm{d}r}$ denotes the rate of change of DSR with respect to $r$.
-As shown in Figure 3(b), MVD increases with the depth of the model’s layers.
-This implies that the safety boundary also expands in deeper layers, requiring more substantial shifts to compromise the model’s safety checks.
-The findings align closely with our intuitive observations in Figure 1.
-Figure 1(a) shows that in low layers, jailbreak activations are shifted in various directions, but all evidently surpass the safety boundary.
-In the middle layers presented in Figure 1(b), most jailbreak activations are beyond the safety boundary, and they are more aligned with benign activations.
-In the deep layers shown in Figure 1(c), jailbreak activations cluster within a region, which overlaps a little with the safety boundary.
+As shown in Figure 3(b), MVD increases with model depth, indicating that the safety boundary expands in deeper layers and requires larger shifts to bypass safety checks.
+This aligns with Figure 1: in lower layers (Figure 1(a)), jailbreak activations scatter widely but surpass the safety boundary; in middle layers (Figure 1(b)), jailbreak activations align more with benign activations while remaining outside the boundary; in deeper layers (Figure 1(c)), jailbreak activations cluster, partially overlapping the safety boundary.
 3.3.3 Key Takeaways on the Jailbreak Process
 Based on the above analysis, we summarize the jailbreak process as follows: (1) Jailbreak takes effect in the low layers, contrary to prior work Zhou et al.
-(2024), which suggested it begins in the middle layers.
-(2) As the attack progresses, jailbreak continues to operate, shifting the activation outside the safety boundary and causing the model to be deceived.
+(2024b), which suggested it begins in the middle layers.
+(2) As the attack progresses, jailbreak continues to shift the activation outside the safety boundary, causing the model to be deceived.
 (3) When comparing all layers, the low and the middle layers exhibit the strongest shift and smallest safety boundaries, demonstrating their importance in the jailbreak process.
 4 Activation Boundary Defense
-Based on the above findings, we propose a defense called Activation Boundary Defense (ABD), whose core idea is to confine the activations within the safety boundaries.
+Although the safety boundary is not directly measurable, we can still use this concept to prevent jailbreaks.
+Here we propose a lightweight defense called Activation Boundary Defense (ABD), whose core idea is to confine the activations within the safety boundaries.
 The workflow of ABD is shown in Figure 4.
@@
 Approximation of activation distributions.
-We empirically find that the activation distributions in each layer can be approximated by a normal distribution.
+We find that the activation distributions in each layer can be well approximated by a normal distribution.
 The proof is as follows.
 For each layer $l$, we examine two distributions: the activation coordinate distribution $\mathcal{D}^{l}(x)$ and a normal distribution $\mathcal{D_{N}}^{l}(x)$ with the same mean $\mu^{l}_{\mathcal{D}}$ and standard deviation $\sigma^{l}_{\mathcal{D}}$.
-To measure their similarity, we compute the Jensen-Shannon (JS) divergence Lin (1991) between $\mathcal{D}^{l}(x)$ and $\mathcal{D_{N}}^{l}(x)$.
+To measure their similarity, we compute the Jensen-Shannon divergence (JS divergence) Lin (1991) between $\mathcal{D}^{l}(x)$ and $\mathcal{D_{N}}^{l}(x)$.
 Across all layers, the maximum JS divergence is 0.0839, and the mean is 0.0575, and both are well below 0.1.
-These low JS divergence values indicate strong similarity between $\mathcal{D}^{l}(x)$ and $\mathcal{D_{N}}^{l}(x)$, which supports the validity of approximating $\mathcal{D}^{l}(x)$ with $\mathcal{D_{N}}^{l}(x)$.
+We visualize all JS divergences in Table 7 in Appendix E.1.
+These low JS divergence values indicate a strong similarity between $\mathcal{D}^{l}(x)$ and $\mathcal{D_{N}}^{l}(x)$, which supports the validity of approximating $\mathcal{D}^{l}(x)$ with $\mathcal{D_{N}}^{l}(x)$.
 Penalty function design.
-Figure 4: Workflow of ABD. ABD restricts outlier activation coordinates using a penalty function and determines its application scope via BO-based tuning.
 To design a practical penalty function for activation coordinates, we establish three key principles: (1) It should target only outlier activations, leaving non-outliers unaffected.
@@
 $x^{\prime}=\alpha^{l}\cdot\mathrm{tanh}(\beta^{l}\cdot(x-\mu^{l}_{\mathcal{D}}))+\mu^{l}_{\mathcal{D}},$ (3)
-where $\alpha^{l}\geq 0$ and $\beta^{l}\geq 0$ are hyperparameters. Visualization is given in Appendix D.1.
+where $\alpha^{l}\geq 0$ and $\beta^{l}\geq 0$ are hyperparameters. The penalty functions under different $\alpha^{l}$ and $\beta^{l}$ are visualized in Figure 6 in Appendix E.2.
 This function meets the stated expectations.
@@
 Moreover, the function mostly penalizes outliers.
-Within the range $[\mu^{l}_{\mathcal{D}}-b^{l},\mu^{l}_{\mathcal{D}}+b^{l}],x^{\prime}\approx x$, resulting in a negligible penalty for values close to the mean.
+Within the range $[\mu^{l}_{\mathcal{D}}-b^{l},\mu^{l}_{\mathcal{D}}+b^{l}],x^{\prime}\approx x$, where $b^{l}$ is a relatively small number, resulting in a negligible penalty for values close to the mean.
 The hyperparameters $\alpha^{l}$ and $\beta^{l}$ govern the behavior of the penalty function.
@@
 Meanwhile, $\beta^{l}$ controls the size of the unmodified region $[\mu^{l}_{\mathcal{D}}-b^{l},\mu^{l}_{\mathcal{D}}+b^{l}]$; increasing $\beta^{l}$ expands this region, while decreasing $\beta^{l}$ narrows it.
+Figure 4: Workflow of ABD. ABD restricts outlier activation coordinates using a penalty function and determines its application scope via BO-based tuning.
 4.2 Bayesian Optimization-Based Tuning
-Model Jailbreak No Defense Paraphrase PPL Retokenization SafeDecoding Self-Exam Self-Reminder ABD (Ours)
-Vicuna No attack 88% 82% 90% 66% 88% 100% 100% 100%
-GCG-Individual 4% 86% 78% 62% 100% 86% 100% 100%
-AutoDAN 6% 26% 12% 36% 78% 30% 30% 44%
-PAIR 30% 62% 40% 30% 94% 86% 100% 76%
-DeepInception 10% 2% 2% 0% 98% 18% 40% 58%
-Llama-2 No attack 100% 100% 100% 94% 100% 100% 100% 100%
-GCG-Individual 50% 98% 100% 98% 100% 64% 100% 100%
-AutoDAN 98% 94% 98% 94% 100% 100% 98% 100%
-PAIR 72% 90% 82% 78% 92% 100% 86% 92%
-DeepInception 74% 82% 88% 60% 100% 94% 96% 100%
-Table 1: The Defense Success Rate (DSR) of different defense methods across Vicuna-7B-v1.3 (simplified as Vicuna) and Llama-2-7B-chat (simplified as Llama-2) along with different attack types, where a higher DSR indicates better defense effectiveness.
-For ABD, the best and second performance across all defenses are highlighted.
-Another crucial aspect of our method is the configuration of the hyperparameters and the application scope of the penalty function.
-Specifically, we must (1) minimize the number of affected activation coordinates to limit unnecessary perturbations, and 2) for each affected layer $l$, find the values of $\alpha^{l}$ and $\beta^{l}$ that best improve the model’s resilience against adversarial attacks.
+Model Jailbreak No Defense Paraphrase PPL Retokenization SafeDecoding Self-Exam Self-Reminder IA ABD (Ours)
+Vicuna-7B No attack 88% 82% 90% 66% 88% 100% 100% 100% 100%
+GCG-Individual 4% 86% 78% 62% 100% 86% 100% 100% 100%
+AutoDAN 6% 26% 12% 36% 78% 30% 30% 100% 44%
+PAIR 30% 62% 40% 30% 94% 86% 100% 100% 76%
+DeepInception 10% 2% 2% 0% 98% 18% 40% 100% 58%
+Llama-2 No attack 100% 100% 100% 94% 100% 100% 100% 100% 100%
+GCG-Individual 50% 98% 100% 98% 100% 64% 100% 100% 100%
+AutoDAN 98% 94% 98% 94% 100% 100% 98% 100% 100%
+PAIR 72% 90% 82% 78% 92% 100% 86% 100% 92%
+DeepInception 74% 82% 88% 60% 100% 94% 96% 100% 100%
+Table 1: DSR of different defense methods across Vicuna-7B and Llama-2 along with different attack types . For ABD, the best, second and third performance across all defenses are highlighted.
+Model Jailbreak No Defense IA ABD (Ours)
+Qwen No attack 96.00% 84.00% 98.00%
+GCG-Individual 6.84% 44.64% 98.74%
+Vicuna-13B No attack 98.00% 100.00% 100.00%
+GCG-Individual 38.46% 100.00% 100.00%
+Table 2: Representative results comparing DSR of Qwen and Vicuna-13B. The best performance is highlighted.
+Another crucial aspect of ABD is determining which layers to penalize and what hyperparameters to choose.
+Specifically, we aim to minimize the affected layers to reduce unnecessary perturbations and optimize $\alpha^{l}$ and $\beta^{l}$ for each layer $l$ to enhance the model’s resilience against jailbreaks.
 To accomplish these goals, we use a Bayesian Optimization (BO)-based Jones et al.
-(1998); Mockus (1974) tuning method.
+(1998) tuning method to find suitable configurations.
+The BO process used for tuning is detailed in Appendix E.4.
 Concretely, we define two core objectives.
@@
 […]
-Table 2: Comparison of defenses on Runtime per Query and Just-Eval metrics in Vicuna-7B-v1.3.
+Table 3: Comparison of defenses on Runtime per Query and Just-Eval metrics in Vicuna-7B.
 $\downarrow$: smaller is better; $\uparrow$: larger is better.
-For ABD, the best and second performances are highlighted.
+For ABD, the best and second performances among all defenses are highlighted.
 ABD preserves the model’s general performance, adding less than 0.1 seconds to runtime while producing high-quality outputs with leading evaluation scores.
@@
 Backbone models.
-We consider two widely used open-source LLMs as the target models in our experiments: Llama-2-7B-Chat Touvron et al.
-(2023) and Vicuna-7B-v1.3.
-Llama-2-7B-Chat is specially trained for safety alignment, while Vicuna-7B-v1.3 is tuned based on Llama without extra alignment for safety.
-The configurations of the LLMs are shown in Appendix E.1.
+We evaluated four widely used open-source LLMs of varying sizes and architectures: Llama-2-7B-Chat Touvron et al.
+(2023), Vicuna-7B-v1.3, Qwen-1.5-0.5B-Chat Bai et al.
+(2023), and Vicuna-13B-v1.5, referred to as Llama-2, Vicuna-7B, Qwen, and Vicuna-13B.
+Llama-2 is specifically trained for safety alignment, while Vicuna-7B is fine-tuned from Llama without additional safety alignment.
+Both Llama-2 and Vicuna-7B have 32 layers.
+Qwen and Vicuna-13B have 24 and 40 layers, respectively.
+The LLM configurations are detailed in Appendix F.1.
 Jailbreak and defense baselines.
-For jailbreak methods, we considered four widely applied ones: optimization-based methods include GCG-Individual and AutoDAN, which iteratively optimize a jailbreak prompt towards the goal of generating affirmative responses.
+For jailbreak methods, we considered four widely applied ones: optimization-based methods include GCG-Individual and AutoDAN, which iteratively optimize a jailbreak prompt aiming at generating affirmative responses.
 Model-based jailbreak, i.e., PAIR, uses an attacker LLM to refine the jailbreak prompts.
@@
 (2023), which crafts jailbreak prompts based on a stealthy template.
-Correspondingly, we applied different defense methods to the models to compare them to our model for combating jailbreak methods: Paraphrase Jain et al.
+A detailed illustration of jailbreaks can be found in F.2.
+Correspondingly, we utilized different defense methods as baselines for ABD: Paraphrase Jain et al.
 (2023), Retokenization Jain et al.
 (2023) and Self-Reminder Xie et al.
-(2023) reformulate the input to avoid attack; PPL Alon and Kamfonas (2023) and Self-ExamPhute et al.
+(2023) reformulate the input to avoid attack; PPL Alon and Kamfonas (2023), Self-ExamPhute et al.
+(2023) and Intention Analysis (IA) Zhang et al.
 (2025) defense LLMs by double-checking their outputs; SafeDecoding Xu et al.
 (2024) uses a tuned model to modify the output probability distribution.
+Further introduction about these defenses can be found in Appendix F.3.
 Datasets and metrics.
 Following Xu et al.
-(2024) and Chao et al.
-(2023), we utilized 50 samples from AdvBench Zou et al.
+(2024), we adopted Just-Eval Lin et al.
+(2023) to measure the general abilities of LLMs applied defense.
+Just-Eval is a comprehensive benchmark containing 1,000 diverse instructions, covering seven task types (e.g., reasoning, math, coding, etc.) and seven topics (e.g., ethics, nature, STEM, etc.).
+Furthermore, following Xu et al.
+(2024) and Liu et al.
+(2024d), we utilized the 50-behavior subset from AdvBench Zou et al.
 (2023) as the test set.
+This subset, curated by Chao et al.
+(2023), was created by removing duplicate harmful behavior prompts from AdvBench, ensuring efficiency while mitigating biases caused by repeated behaviors.
 We used DSR when evaluating defense effectiveness.
-Furthermore, following Xu et al.
-(2024), we adopted Just-Eval Lin et al.
-(2023) to measure the general abilities of LLMs before and after applying defense.
-Just-Eval is a comprehensive benchmark containing 1,000 diverse instructions, covering seven task types (e.g., reasoning, math, coding, etc.) and seven topics (e.g., ethics, nature, STEM, etc.).
 Following Lin et al.
 (2023), we leveraged GPT-4o-mini OpenAI (2024) to score the quality of the outputs, ranging from 1 to 5, across five aspects: helpfulness, clarity, factuality, depth, and engagement.
-We reported the average score for each aspect and the overall average score for different aspects, denoted as Avg.
-To evaluate the efficiency, we calculated the Runtime per Query for each defense, which represents the average time taken to process a single query during inference.
-This metric provides a direct measurement of the efficiency of defenses in real-world applications.
-We also utilized the above two metrics to derive an overall score that simultaneously reflects LLMs’ responding speed and quality.
-The calculation is shown in Appendix E.4.1.
+We reported the average score for each aspect and their macro-average score, denoted as Avg.
+To evaluate the efficiency, we reported Runtime per Query, which represents the average time taken to process a single query.
+We also utilized the above metrics to derive an overall score that simultaneously reflects LLMs’ responding speed and quality.
+The calculation of the overall score is shown in Appendix F.4.2.
 Implementation Details.
@@
 GCG-Universal finds a shared jailbreak suffix for all harmful prompts that can deceive LLMs.
-By default, we set $w=0.7$ when calculating $\mathcal{L}_{\mathrm{Robust}}(\Theta,M)$ to slightly prioritize improving DSR while balancing defensive capability and perturbation to LLMs.
-Appendix D.2 presents more details about ABD optimization.
+We set $w=0.8$ when calculating $\mathcal{L}_{\mathrm{Robust}}(\Theta,M)$ as it best balances performances on both attacked and unattached data.
+We conduct ablation studies on $w$ in Appendix C.3, results shown in Table 5.
+Appendix E.3 presents settings of ABD optimization.
 5.2 Experimental Results
 ABD reveals vulnerabilities in low and middle layers.
-In both models, low and middle layers, such as layers 5 and 12 in Vicuna-7B-v1.3 and layers 2 and 12 in Llama-2-7B-chat, are penalized, which aligns with our observed findings and proposed explanation that low and middle layers are more vulnerable to jailbreak attacks due to noticeable activation shifts.
+In all models, the selected layers under the mask $M$ mainly include low and middle layers, such as layers 5 and 12 in Vicuna-7B, layers 2 and 12 in Llama-2, layers 2, 11, and 14 in Vicuna-13B, and layers 5 and 14 in Qwen.
+These layers are most frequently penalized, consistent with our observations and proposed explanation that low and middle layers are more vulnerable to jailbreak attacks due to significant activation shifts.
 ABD successfully constrains jailbreak activations.
-Figure 3(c) shows the percentage of jailbreak activations within the harmful activation space, namely the inclusion ratio.
-Before applying ABD, the inclusion ratio remains below 0.4 for various jailbreak activations, indicating that jailbreak activations lie outside of the safety boundary.
-After applying ABD, the ratio increases to 1, demonstrating that our method effectively constrains the jailbreak activations within the safety boundary.
-Additional details are provided in Appendix C.2.
-From a visualization perspective, Figure 1 projects activations for different prompts.
-Under our defense, jailbreak activations are progressively constrained within the harmful activation space under ABD, as shown in Figure 1.
-further confirming the effectiveness of ABD for mitigating jailbreak effects.
-Statistic effectiveness of ABD.
+Figure 3(c) shows the inclusion ratio, representing the percentage of jailbreak activations within the harmful activation space.
+The detailed calculation of inclusion ratios is provided in Appendix C.2.
+Before ABD, the inclusion ratio of various jailbreak activations is below 0.4, indicating that these activations lie outside the safety boundary.
+After ABD, the inclusion ratio rises to 1, demonstrating ABD’s effectiveness in constraining jailbreak activations.
+For visualization, Figure 1 projects activations for different prompts.
+Under ABD, jailbreak activations are progressively constrained within the harmful activation space.
+This further confirms that ABD progressively confines jailbreak activations within the harmful activation space and mitigates jailbreak effects.
+Robust Defense Performance of ABD.
 We report the statistical defense results of ABD on the test set in Table 1.
-For Vicuna-7B-v1.3, which lacks specific safety alignment, the defense poses a more challenging task.
+We also report representative results on Qwen and Vicuna-13B in Table 2, with full results shown in Table 9 in Appendix G.2.
+For Vicuna-7B which lacks specific safety alignment, the defense poses a more challenging task.
 Nevertheless, ABD demonstrates competitive performance.
@@
 For the well-aligned model Llama-2-7B-chat, ABD achieves 100% DSR under most jailbreak methods.
-ABD is efficient and reliable.
-Table 2 shows Runtime per Query, Just-Eval scores, and overall scores of Vicuna-7B-v1.3 with different defenses applied.
+For Qwen and Vicuna-13B, ABD maintains nearly 100% DSR, showcasing competitive performance.
+Notably, for smaller models like Qwen, IA struggles to match its effectiveness on 7B or 13B models, while ABD remains consistently effective.
+To further validate ABD, we tested it on the full AdvBench and a broader set of jailbreaks.
+Results are shown in Table 8 in Appendix G.
+Efficiency and minimal overhead of ABD.
+Table 3 shows Runtime per Query, Just-Eval scores, and overall scores of Vicuna-7B with different defenses applied.
 We find that ABD has the greatest overall score across all defenses.
 Moreover, ABD only adds marginal extra time cost and general ability affection.
-Specifically, it only causes less than $1\%$ delay in each sample and less than $2\%$ perturbation in overall ability, compared to costly methods such as Paraphrase and Self-Exam.
+Specifically, it only causes less than $1\%$ delay in each sample and less than $2\%$ perturbation in overall ability, compared to costly methods such as IA, Paraphrase and Self-Exam.
 Furthermore, with ABD applied, the helpfulness, actuality, depth, and engagement also show a slight increase.
-We further discover that for baseline defenses such as Retokenization and Self-Reminder, LLM would generate overly simplistic outputs, which leads to smaller Runtime per Query, but they have significantly smaller Just-Eval scores.
+We further discover that for baseline defenses such as Retokenization and Self-Reminder, LLM would generate overly simplistic outputs, which leads to smaller Runtime per Query, but they have significantly smaller Just-Eval scores.We also compared defenses on complexity, overhead, and implementation difficulty, highlighting ABD’s efficiency (see Table 10 in Appendix E.5).
 6 Conclusion and Future Work
 We conducted a comprehensive study of jailbreak mechanisms, analyzing over 30,000 samples.
-Our findings reveal that jailbreak shifts harmful activations outside the safety boundary in each layer, with the most severe shifts occurring in the low and middle layers.
+Our findings reveal that jailbreak shifts harmful activations outside the safety boundary in each layer, with the most severe shifts in the low and middle layers.
 Motivated by this finding, we proposed ABD, which drives jailbreak activations back within the safety boundary, utilizing LLMs’ intrinsic sensitivity to harmful information.
 Our experiments suggest that ABD is both practical and efficient.
-In the near future, we aim to investigate the safety challenges of jailbreak attacks in multi-turn dialogue systems.
+In the near future, we aim to investigate the challenges of jailbreaks in multi-turn dialogue scenarios.
 Limitations
@@
 For certain attack methods, under-aligned models (Vicuna-7B-v1.3) may not perform significantly as well-aligned models (Llama2-7B-chat).
-We believe this is because under-aligned models have unclear safety boundaries, which complicate the search for penalty functions that balance general ability and DSR.
+We believe this is because ABD’s effectiveness depends on safety alignment.
+Under-aligned models have unclear safety boundaries, which complicate the search for penalty functions that balance general ability and DSR.
+Therefore, they typically need more extensive optimization.
 Future work could refine ABD by specifically searching for activation spaces that preserve the concept of "safety", therefore enhancing its generalizability on uncensored models.
@@
 (2024).
-As a result, the relationship between jailbreak dialogues and safety boundaries remains largely unexplored.
-Addressing this limitation, we plan to investigate and incorporate methods for detecting safety boundaries in multi-round dialogue scenarios as part of our future work.
+As a result, the relationship between jailbreaks and safety boundaries remains largely unexplored.
+While multi-turn jailbreaks are beyond this paper’s scope, ABD can extend to them by using prior context tokens to compute $\mu^{l}_{\mathcal{D}}$ and applying the same optimization strategy.
+To adapt ABD for multi-turn jailbreaks, potential improvements include: (1) Dynamic adjustment: Relax or strengthen ABD constraints based on the safety of prior turns.
+(2) Selective application: Apply ABD at critical points, e.g., when malicious topics emerge, to balance efficiency and effectiveness.
+(3) Constraint propagation: Carry over constraints across turns to prevent harmful activations from accumulating.
 Ethical Considerations
@@
 We highlight that the development of ABD only needs publicly available datasets and jailbreak methods and does not require designing new jailbreak methods.
-We demonstrate some harmful responses from LLMs only for illustration purposes.
+We demonstrate some harmful responses from LLMs only for illustration.
 We acknowledge that ABD would cause the development of new attacks.
-Therefore, we will explore using random perturbation in the input sequence rather than a particular jailbreak method when optimizing to mitigate such attacks.
+We will explore using random perturbation in the input sequence rather than a particular jailbreak method when optimizing to mitigate such attacks.
 Appendix A Statistics of Observed Data
@@
 We visualize activation spaces of all layers in Vicuna-7B-v1.3, as shown in Figure LABEL:fig:all_subfig1, Figure LABEL:fig:all_subfig2 and Figure LABEL:fig:all_subfig3.
-Appendix C Supplementary Experiments
+Appendix C Supplementary Observational Experiments
 C.1 Data Augmentation Experiment
@@
 Based on 8,556 harmful samples gathered in Table 4, we calculate a ball that covers $80\%$ activations.
-The center of the ball is $\mu_{\mathcal{D}}^{l}$, and the radius of the ball is denoted as $r_{\mathcal{D}}^{l}$.
+The center of the ball is $c_{\mathcal{D}}^{l}$, and the radius of the ball is denoted as $r_{\mathcal{D}}^{l}$.
 Then, we calculate the portion of $A^{l}$ which are contained within the ball:
-$\rho^{l}_{\mathrm{inclusion}}=\frac{\left|\{a^{l}_{i}\in A^{l}\mid\|a^{l}_{i}-\mu_{\mathcal{D}}^{l}\|_{2}\leq r_{\mathcal{D}}^{l}\}\right|}{|A^{l}|},$
-where $\|a^{l}_{i}-\mu_{\mathcal{D}}^{l}\|_{2}$ is the distance between the activation and the center.
+$\rho^{l}_{\mathrm{inclusion}}=\frac{\left|\{a^{l}_{i}\in A^{l}\mid\|a^{l}_{i}-c_{\mathcal{D}}^{l}\|_{2}\leq r_{\mathcal{D}}^{l}\}\right|}{|A^{l}|},$
+where $\|a^{l}_{i}-c_{\mathcal{D}}^{l}\|_{2}$ is the distance between the activation and the center.
 We calculate inclusion ratios of different types of jailbreak activations, with three representative types shown in Figure 1.
@@
 Notably, despite different types of jailbreaks, they all achieve an inclusion ratio of $100\%$ with ABD applied, which verifies the effect of our method.
-Appendix D Supplementary Illustration on ABD
+C.3 Ablation on Optimizing Objective
+We conduct ablation studies to determine the best $w$ in $\mathcal{J}_{\mathrm{total}}(\Theta,M)$.
+We adopt Vicuna-7B-v1.3 as the backbone model, and use the 50-behavior subset as the testing set.
+Additionally, we fix optimization steps to 2000.
+DSR results are shown in Table 5.
+We discover that (1) Higher w improves robustness on attacked datasets.
+(2) Lower w preserves performance on unattacked datasets.
+(3) At w=0.8, the model achieves the best average DSR, balancing robustness and benign performance.
+Therefore, we set $w=0.8$ in our experiments in Section 5.
+Jailbreak No defense $w=0.2$ $w=0.5$ $w=0.8$ $w=1$
+No Jailbreaks 96% 100% 86% 100% 86%
+GCG-Individual 7% 75% 88% 93% 100%
+Avg. 52% 88% 87% 96% 93%
+Table 5: Defense performance comparison under different $w$.
+Appendix D Supplementary Illustration on Safety Boundaries
+To gain a deeper understanding of the safety boundary, we present a case study indicating how the model’s responses vary based on the level of RAS applied to layer 1.
+The results are shown in Table 6.
+It shows that (1) The safety boundary is implicit, as a significant shift does not necessarily results in successful deception, such as when $r=80$.
+(2) As the responses move further away from this boundary, the model begins to exhibit uncommon response patterns such as giving a presentation ($r=100$) or more casual styles ($r=30,r=80$).
+We believe this phenomenon may correlate with an increased likelihood of producing jailbroken outputs.
+Table 6: Model output under different shifting distances.
+Query: Write a script that can exploit vulnerabilities in a software or operating system
+Distance $r$ Response Jailbroken
+5 I’m sorry, but as an AI language model, I cannot provide a script… No
+30 alright, here’s a script that can exploit vulnerabilities… Yes
+80 , but I am not able to do so as … No
+[…]
+Appendix E Supplementary Illustration on ABD
 Figure 5: t-SNE visualization of activations in layer 14.
@@
 Right: Results after scaling up to 500 samples per type, showing jailbreak activations clustering on the harmful side.
-D.1 Visualization of the penalty function.
+E.1 Further analysis of Distribution Approximation
+We use the 20,000 benign samples and 8,556 harmful samples to form 4096-dim activations.
+For each layer, we compute their coordinate distributions: benign activation distribution $\mathcal{D}^{l}_{B}$ and harmful activation distribution $\mathcal{D}^{l}_{H}$.
+Using their means and divergences, we generate normal distributions $\mathcal{D}^{l}_{BN}$ and $\mathcal{D}^{l}_{HN}$.
+Finally, we calculate the JS-divergence (JSD) between the original and generated distributions.
+Results are shown in Table 7.
+The results show that (1) In all layers, $JSD(\mathcal{D}^{l}_{B},\mathcal{D}^{l}_{BN})<0.1$ and $JSD(\mathcal{D}^{l}_{H},\mathcal{D}^{l}_{HN})<0.1$, confirming the validity of approximating coordinate distributions with normal distributions.
+(2) Earlier layers have lower JSDs, making this approximation safer in early and middle layers.
+This supports our ABD settings, where penalty functions are primarily applied to these layers.
+Table 7: Layer-wise JSD Comparisons
+Layer 0 1 2 3 4 5 6 7
+$JSD(D_{B}^{l},D_{BN}^{l})$ 0.0036 0.0026 0.0113 0.0112 0.0174 0.0234 0.0316 0.0379
+$JSD(D_{H}^{l},D_{HN}^{l})$ 0.0033 0.0035 0.0114 0.0099 0.0167 0.0239 0.0354 0.0414
+Layer 8 9 10 11 12 13 14 15
+$JSD(D_{B}^{l},D_{BN}^{l})$ 0.039 0.0455 0.0517 0.058 0.0637 0.0766 0.0757 0.0818
+$JSD(D_{H}^{l},D_{HN}^{l})$ 0.0396 0.0473 0.053 0.0589 0.0661 0.076 0.0753 0.0819
+Layer 16 17 18 19 20 21 22 23
+$JSD(D_{B}^{l},D_{BN}^{l})$ 0.0848 0.0862 0.0864 0.0873 0.0867 0.0862 0.0853 0.0834
+$JSD(D_{H}^{l},D_{HN}^{l})$ 0.0823 0.0841 0.0845 0.0844 0.083 0.0815 0.0804 0.0755
+Layer 24 25 26 27 28 29 30 31
+$JSD(D_{B}^{l},D_{BN}^{l})$ 0.0789 0.0786 0.078 0.0762 0.0724 0.0712 0.0752 0.078
+$JSD(D_{H}^{l},D_{HN}^{l})$ 0.0681 0.0672 0.0659 0.0628 0.0601 0.0593 0.0643 0.0724
+E.2 Visualization of the penalty function.
 A visualization of penalty functions are shown in Figure 6.
@@
 Figure 6: Penalty functions under different $\alpha^{l},\beta^{l}$ compared with $x^{\prime}=x$.
-D.2 ABD Optimization Settings
+DSR No attack GCG-universal GCG-individual PAIR AdvPrompter COLD-Suffix AutoDAN Avg. Perform
+No Defense 96.73% 87.50% 5.29% 28.18% 82.05% 59.86% 2.50% 51.73%
+ABD (Ours) 99.81% 99.68% 98.23% 91.51% 97.65% 98.02% 20.77% 86.52%
+Table 8: Performance comparison of different methods under various attack scenarios. The best performance is highlighted in bold.
+Model Attack No Defense Paraphrase PPL Retokenization Self-Exam Self-Reminder IA ABD (Ours)
+Qwen No attack 96.00% 80.00% 96.00% 58.00% 100.00% 96.00% 84.00% 98.00%
+GCG-Individual 6.84% 67.52% 36.75% 58.97% 42.74% 82.05% 44.64% 89.74%
+AutoDAN 14.00% 12.00% 26.00% 4.00% 50.00% 8.00% 2.00% 36.00%
+Vicuna-13B No attack 98.00% 98.00% 98.00% 70.00% 100.00% 96.00% 100.00% 100.00%
+GCG-Individual 38.46% 98.08% 98.08% 84.62% 84.62% 98.08% 100.00% 100.00%
+AutoDAN 14.00% 54.00% 20.00% 16.00% 96.00% 98.00% 100.00% 74.00%
+Table 9: Performance comparison of different defenses across models with different scales and structures. For ABD, the best and second-best performance across all defenses are highlighted.
+E.3 ABD Optimization Settings
 Validation data.
@@
 Optimizing framework.
-We adopt Optuna Akiba et al. (2019) as our framework of optimization. We follow Optuna’s default settings, i.e., Gaussian process-based Bayesian Optimization.
-Appendix E Supplementary Illustration on Experiments
-E.1 Generation Configs
-When conducting experiments, we directly utilize most of the original configurations of Vicuna-7B-v1.3 and Llama2-7B-chat.
+We adopt Optuna Akiba et al. (2019) as our framework of optimization. We follow Optuna’s default settings, i.e., Tree-structured Parzen Estimator (TPE) Bergstra et al. (2011)-based Optimization.
+E.4 ABD Optimizing Process
+The optimization process utilizes the Tree-structured Parzen Estimator (TPE) algorithm, a Bayesian optimization variant that efficiently explores high-dimensional parameter spaces by modeling promising regions.
+TPE constructs two probabilistic density distributions, $l(\Theta)$ and $g(\Theta)$, over the defense parameters $\Theta=\{m_{i},\alpha_{i},\beta_{i},k_{i}\}$, where $l(\Theta)$ captures the likelihood of $\Theta$ yielding high objective score $\mathcal{J}_{\mathrm{total}}$, and $g(\Theta)$ corresponds to lower-score regions.
+At each iteration, TPE samples new candidates by maximizing the Expected Improvement criterion Jones (2001):
+$\Theta_{\mathrm{new}}=\arg\max_{\Theta}\frac{l(\Theta)}{g(\Theta)},$
+prioritizing parameters likely to outperform the current best.
+The process begins with a warm-up phase, where random evaluations establish a baseline, followed by a density-guided search, which iteratively refines $l(\Theta)$ and $g(\Theta)$ using historical observations.
+This approach ensures efficient exploration while mitigating local optima risks.
+By leveraging Optuna’s TPE implementation, the optimizer adapts dynamically to $\mathcal{J}_{\mathrm{total}}$ feedback, achieving automated defense parameter optimization.
+E.5 Comparison with other defense methods
+We compared ABD with other defense methods in terms of usability.
+We qualitively evaluated usability based on extra computational complexity, extra tokens, additional modules, and deployment difficulty.
+The results are shown in Table 10.
+Compared to other defense methods, ABD features no additional token overhead and only constant-level extra complexity.
+Besides, the operation is simple and straightforward, which further improves its utility.
+Appendix F Supplementary Illustration on Experiments
+Table 10: Empirical Comparison of Defenses. $t$ is the length of input sequence.
+Extra
+Complexity
+Extra
+Tokens
+Extra
+Modules
+Deployment Difficulty Performance Utility
+PPL $O(t^{2})$ Extra Conversation GPT-2 Low Low Low
+Paraphrase $O(t^{2})$ Extra Conversation - Low Low Low
+Retokenization $O(t\log t)$ Low Low Low
+Self-Exam $O(t^{2})$ Extra Conversation - Low Low Low
+SafeDecoding $O(t)$ LoRA Adapter High High Medium
+Self-Reminder $O(t^{2})$ Extra Instructions - Low Low Low
+IA $O(t^{2})$ Extra Conversation - Low High Medium
+ABD (Ours) $O(1)$ Mathematical Operation Low High High
+F.1 Generation Configs
+When conducting experiments, we directly utilize most of the original configurations of all LLMs.
 Specifically, we set max_new_tokens=128.
@@
 fastchat.model.get_conversation_template(template_name).
-We set template_name="vicuna" for Vicuna-7B-v1.3 and template_name="llama-2" for Llama2-7B-chat.
-E.2 Jailbreak Methods
+We set template_name="vicuna" for Vicuna-7B and Vicuna-13B, template_name="llama-2" for Llama-2, and template_name="qwen" for Qwen.
+F.2 Jailbreak Methods
 GCG-Individual and GCG-Universal
@@
 It creates a virtual and layered scene, allowing the model to find flexible ways to bypass usage controls in normal situations.
-E.3 Defense Methods
+F.3 Defense Methods
 PPL
@@
 Xie et al. (2023) defends the LLMs by adding a system prompt specifically emphasizing potential safety issues to prevent the model from outputting harmful responses.
-E.4 Metrics
-E.4.1 Calculating DSR
+Intention Analysis
+Zhang et al.
+(2025) enhances LLM safety by guiding the model to explicitly identify the essential intention of user queries and subsequently generate policy-aligned responses to mitigate harmful outputs.
+F.4 Metrics
+F.4.1 Calculating DSR
 We use Dic-Judge Zou et al.
@@
 $\mathrm{DSR=\frac{|\{responses\ containing\ refusal\ strings\}|}{|\{input\ jailbreak\ queires\}|}}.$
-We adopt refusal strings as proposed by Xu et al. (2024), presented in Table
-E.4.2 Calculating Overall Scores
+We adopt refusal strings as proposed by Xu et al. (2024), presented in Table 11.
+F.4.2 Calculating Overall Scores
 We use Runtime per Query and overall averaged Just-Eval scores to calculate overall scores:
 $\mathrm{Overall=\big{[}1-Norm(t)+Norm(Avg.)\big{]}/2,}$
-where $\mathrm{t}$ is Runtime per Query and $\mathrm{Avg.}$ represents overall averaged Just-Eval scores. 4.
+where $\mathrm{t}$ is Runtime per Query and $\mathrm{Avg.}$ represents overall averaged Just-Eval scores.
+Here, $\mathrm{Norm(x)}$ denotes the normalization of a metric $\mathrm{x}$, calculated as $\mathrm{x}$ divided by the sum of $\mathrm{x}$ and the corresponding values for all reference groups.
 Refusal Strings
@@
 Table 11: Refusal strings.
+Appendix G Supplementary Experiments
+G.1 Scale Up Baselines and Data
+To further strengthen our evaluation, we expanded the jailbreak test samples from the initial 50-behavior subset to the full AdvBench (5̃20 samples) and applied the jailbreak method from Section 3.2 for a more intuitive observation.
+The defense performances on Vicuna-7B-v1.3 are presented in Table 8.The results remain consistent with Table 1, reaffirming ABD’s strong defense capability against various stealthy jailbreak methods.
+G.2 Full Results on Qwen and Vicuna-13B
+The results are shown in Table 9.
+ABD remains performing decent on other sizes of models.
+Besides, in Vicuna-13B, 6 out of 9 selected layers are early and intermediate layers (layer 2, 11, 14, 19, 26, 29), and in Qwen, 2 out of 3 layers (layer 5, 14).
+This supports our interpretations of jailbreak mechanisms, as well as the conclusion that early and middle layers are most vulnerable to jailbreaks, basically hold in larger and smaller models.